### v1

# A decoder-only foundation model for time-series forecasting

Abstract
Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Our model is based on pretraining a patched-decoder style attention model on a large timeseries corpus, and can work well across different forecasting history lengths, prediction lengths and temporal granularities. Author names are listed in alphabetical order.
1 Introduction
Time-series data is ubiquitous in various domains such as retail, finance, manufacturing, healthcare and natural sciences. In many of these domains, one of the most important use-cases of time-series data is forecasting. Time series forecasting is critical to several scientific and industrial applications, like retail supply chain optimization, energy and traffic prediction, and weather forecasting. In recent times, Deep learning models [17, 2] have emerged as a popular approach for forecasting rich, multivariate, time series data, often outperforming classical statistical approaches such as ARIMA or GARCH [3]. In several forecasting competitions such as the M5 competition [12] and IARAI Traffic4cast contest [8], almost all the winning solutions are based on deep neural networks.
At the same time, we are witnessing a rapid progress in the Natural Language Processing (NLP) domain on large foundation models for downstream NLP tasks. Large language models (LLMs) are growing in popularity because they can be used to generate text, translate languages, write different kinds of creative content, and answer your questions in an informative way [16]. They are trained on massive amounts of data, which allows them to learn the patterns of human language. This makes them very powerful tools that can be used for a variety of downstream tasks, often in a zero-shot learning mode.
This motivates the question: “Can large pretrained models trained on massive amounts of time-series data learn temporal patterns that can be useful for time-series forecasting on previously unseen datasets?” In particular, can we design a time series foundation model that obtains good zero-shot out-of-the-box forecasting performance on previously-unseen datasets? Such a time series foundation model, if possible, would bring significant benefits for downstream forecasting users in terms of significantly reduced training data and compute requirements. It is not immediately obvious that such a foundation model for time series forecasting is possible. Unlike in NLP, there is no well defined vocabulary or grammar for time-series. Additionally, the model would need to support forecasting with varying history lengths (context) , prediction lengths (horizon) and time granularities. Furthermore, unlike the huge volume of public text data for pretraining language models, vast amounts of time series data is not readily available. In spite of these issues, we provide evidence to answer the above question in the affirmative.
In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets with different temporal granularities, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets). Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time. The key elements of our foundation model are twofold: 1) a time series corpus built mostly using Google Trends [^1] https://trends.google.com, that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus. Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (225M parameters) and pretraining data size (1B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data.
2 Related Work
In the last decade, deep learning models [17, 2] have emerged as powerful contenders in forecasting time-series in the presence of large training datasets and have been shown to outperform traditional statistical methods such as ARIMA, Exponential smoothing [10]. Forecasting models can be categorized broadly into: (i) Local univariate models that include traditional methods like ARIMA, exponential smoothing [10] and non-autoregressive models like Prophet [19]. These models are trained individually for each time-series in a dataset in order to predict the corresponding time-series’s future. (ii) Global univariate models like DeepAR [17], Temporal Convolutions [2], N-BEATS [14] and long-term forecasting models such as [13, 5] that are trained globally on many time-series but during inference they predict the future of a time-series as a function of its own past and other related covariates. (iii) Global multivariate models that take in the past of all time-series in the dataset to predict the future of all the time-series. Such models include the classical VAR model [27] as well as deep learning models like [18, 25, 28] to name a few.
All the works cited above have primarily been applied in the supervised setting with the notable exception of PatchTST [13] and N-BEATS [14]. PatchTST has a section on dataset to dataset transfer learning in the semi-supervised setting. The patching in our decoder-only model is inspired by [13]. [15] also show that the N-BEATS architecture lends itself to transfer learn between various source-target dataset pairs. However, none of these works aim to train a single foundation model that can work on a plethora of datasets. For a more in-depth discussion about transfer learning in time-series we refer the reader to the survey in [11].
[26] show how to use the the GPT-2 backbone [16] for various tasks including time-series forecasting. [4] is a follow up works along the same lines. Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset. For instance Table-18 [26] shows M4 to M3 transfer. The rest of the two papers are mostly focused on fine-tuning and to the best of our knowledge they do not train a single foundation model that shows out of the box zero-shot performance on a variety of datasets. To the best of our knowledge, the very recent work in TimeGPT-1 [6] is the only known parallel work on foundation model for time-series. However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
3 Problem Definition
The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points. Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy like notation for indices. Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$. Note that since we are building a one-fits-all model we cannot have dataset specific dynamic or static covariates during training time. However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$. See Appendix A.1 for details. Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$. The task is then to learn a capable foundation model that can map any time-series context to horizon, given by
$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L},{\bm{x}}_{1:L+H}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$ (1)
The accuracy of the prediction will be measured by a metric that quantifies their closeness to the actual values. For instance, if the metric is Mean Squared Error (MSE), then the goodness of fit is measured by,
$\mathrm{MSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)=\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}.$ (2)
4 Model Architecture
A foundation model for time-series forecasting should be able to adapt to variable context and horizon lengths, while having enough capacity to encode all patterns from a large pretraining datasets. Transformers have been shown to be able to adapt to different context lengths in NLP [16]. Inspired by the success of patch based modeling in the recent long horizon forecasting work [13] we also chose to breakdown the time-series into patches during training. However, there are several key differences in our foundation model architecture, the primary one being that our model is trained in decoder-only mode [9]. We will now describe the key parts of our architecture and training methodology illustrated in Figure 1.
[…]
[width=]fts_new_v2.pdf
Figure 1: We provide an illustration of our model architecture during training where we show a input time-series of a certain length that can be broken down into input patches. Each patch along with (optional) time-features is processed into a vector by a residual block (as defined in the model definition) to the model dimension of the transformer layers. The vector is then added to positional encodings and fed into $n_{l}$ stacked transformer layers. SA refers to self-attention (note that we use multi-head causal attention) and FFN is the fully connected layer in the transformer. The output tokens are then mapped through a residual block to an output of size output_patch_len which is the forecast for the time-period following the last input patch seen by the model so far.
Input Layers. The job of the input layers is to preprocess the time-series into input tokens to the transformer layers. We first break the input into contiguous non-overlapping patches. Then each patch (along with optional date derived features for that patch) is processed by a Residual Block into a vector of size model_dim. The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [5].
In other words, the inputs $\mathbf{y}_{1:L},{\bm{x}}_{1:L}$ are broken down into patches of size input_patch_len ($p$). The $j$-th patch can be denoted as $\tilde{\mathbf{y}}_{j}=\mathbf{y}_{p(j-1)+1:pj}$ and $\tilde{\mathbf{x}}_{j}=\mathbf{x}_{p(j-1)+1:pj}$. Then the $j$-th input token to the subsequent transformer layers can be denoted as,
$\mathbf{t}_{j}=\texttt{InputResidualBlock}(\tilde{\mathbf{y}}_{j},\tilde{\mathbf{x}}_{j})+\texttt{PE}_{j}$ (3)
where $\texttt{PE}_{j}$ denotes the $j$-th positional encoding as defined in the original transformer paper [20]. There will be $N=\mathopen{}\mathclose{{}\left\lfloor{L/p}}\right\rfloor$ such input tokens.
Stacked Transformer. The bulk of the parameters in our model are in $n_{l}$ transformer layers stacked on top of each other. Each of these layers have the standard multi-head self-attention (MHA) followed by a feed-forward network (FFN). The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads). We set the hidden size of the FFN’s to be equal to model_dim as well. We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token). This can be described by the equation
$\mathbf{o}_{j}=\texttt{StackedTransformer}\mathopen{}\mathclose{{}\left(\mathbf{t}_{1},\cdots,\mathbf{t}_{j}}\right),~{}~{}~{}\forall j\in[N].$ (4)
Output Layers. The remaining task is to map the output tokens into predictions. We train in decoder only mode i.e each output token should be able to be predictive of the part of the time-series that follows the last input patch corresponding to it. This is common for popular large language models like [16]. However, one key difference in our time-series foundation model is that input patch length need not be equal to output patch length i.e we should be able to predict a larger chunk of the time-series based on the encoded information from the input patches seen so far. Let the output patch length be output_patch_len ($h$). We use another Residual Block to map the output tokens to the predictions. This can be described as,
$\hat{\mathbf{y}}_{pj+1:pj+h}=\texttt{OutputResidualBlock}\mathopen{}\mathclose{{}\left(\mathbf{o}_{j}}\right).$ (5)
Thus we encode all the data in $\mathbf{y}_{1:pj}$ into $\mathbf{o}_{j}$ and use that to predict the subsequent $h$ time-points $\mathbf{y}_{pj+1:pj+h}$. This is done for all patches in one training mini-batch.
Loss Function. In this work, we focus on point forecasting. Therefore we can use a point forecasting loss during training like MSE as defined in Equation (2). The loss that is minimized during training can be expressed as,
$\texttt{TrainLoss}=\frac{1}{N}\sum_{j=1}^{N}\mathrm{MSE}\mathopen{}\mathclose{{}\left(\hat{\mathbf{y}}_{pj+1:pj+h},\mathbf{y}_{pj+1:pj+h}}\right).$ (6)
Note that if one is interested in probabilistic forecasting, then it is easy to have multiple output heads for each output patch, each head minimizing a separate quantile loss as in [23]. Another approach can be to output the logits of a probability distribution family and minimize the maximum likelihood loss for probabilistic forecasting [1, 17].
Inference. The trained network can be used to produce forecasts for any horizon using auto-regressive decoding similar to large language models. Given an input $\mathbf{y}_{1:L}$ (assume $L$ is a multiple of $p$ for simplicity) it can first predict $\hat{\mathbf{y}}_{L+1:L+h}$. Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on.
We name our model Pretrained Decoder for Time-series (PreDcT).
5 Empirical Results
We evaluate our model in zero-shot settings on well known public datasets against state-of-the-art supervised forecasting baselines. We show that a single pretrained model can come close or surpass the performance of baselines on the benchmarks even when the baselines are specially trained or tuned for each specific task. Subsequently, we perform ablation studies that justify different choices made in our foundation model architecture.
5.1 Zero-shot Evaluation
We are interested evaluating the performance of a single foundation model pretrained on a large time-series dataset on different target datasets never seen before by the model or, in other words, zero-shot evaluation. To test the generalization performance of our baselines, we choose target datasets covering a variety of data domains, scales and time granularities. We will first discuss the pretraining and target datasets, along with the baselines we use. Then we demonstrate the (Z)ero-(S)hot performance of our model PreDcT(ZS) compared against state-of-the-art models trained directly on the target datasets.
Pretraining Data. We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model. It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model. In this paper, we find that Google Trends [^2] https://trends.google.com can provide a time series corpus that is ideally suited for pre-training our foundation model. Google Trends captures search interest over time for millions of queries. We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022. We use the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset. The date ranges are Jan. 2018 to Dec 2019 for hourly and Jan. 2007 to Dec. 2021 for the other granularities. Along with the trends data, we also add time series from several other publicly available datasets to our pretraining corpus. We add in all the granularities of the M4 dataset [12] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [28]). M4 has a good mix of granularities with around 100k time-series in total. Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points. In addition, we add all the 15 min granularity traffic time series from [22].
We train on a mixture distribution over these datasets that aims to give sufficient weightage to all granularities. We train with a maximum context length of 512 whenever the length of the time-series allows that. For weekly granularity we do not have sufficiently long time-series therefore a max. context length of 256 is used. For the same reason, max. context length of 64 is used while training on monthly and higher granularity data.
Target Datasets. To benchmark our model’s performance, we choose commonly used forecasting datasets of varying sizes that cover various domains, granularities, context lengths and horizon lengths, to test the generalization power of our foundation model against other baselines. The details are summarized in Table 1.
Dataset # Time-Series Time Length Granularity Context Horizon
ETTm1 7 69680 15 Min. 512 96
ETTm2 7 69680 15 Min. 512 96
ETTh1 7 17420 1 Hour 512 96
ETTh2 7 17420 1 Hour 512 96
Wiki 115084 803 1 Day 256 56
ILI 7 966 1 Week 96 24
TourismL 555 228 1 Month 64 12
Table 1: Statistics and task definitions of the target datasets.
(Sub)Hourly.  For 15 min. granularity we use the ETTm1, ETTm2 datasets and test all models on the task of predicting a horizon of 96 time-points after seeing a context of size 512. For hourly granularity, we choose the ETTh1, ETTh2 datasets and test all models on the same task as the 15 min. datasets. The datasets and the context, horizon pair have been widely used in long-term forecasting benchmarks [28, 13]. Note that we used the more challenging original, unscaled versions of these datasets in order to test our model’s zero-shot performance on time-series of different scales.
Daily.  For the daily granularity, we use the Wikipedia web-traffic dataset from the corresponding Kaggle competition  [^3] https://www.kaggle.com/code/muonneutrino/wikipedia-traffic-data-exploration. It has 115k time-series with over two years of data if we exclude the time-series with missing values. The dataset contains web-traffic to Wikipedia articles and the task is to predict the web-traffic (in $\log$ scale) on future dates. We choose a context length of 256 to predict a horizon of 56 days (8 weeks). This dataset has been used in prior multivariate forecasting papers such as like [18].
Weekly.  We use the ILI dataset [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html that collects the number of patients and influenza-like illness ratio in a weekly frequency. We use a context length of 96 to predict 24 weeks into the future. This is one of the configurations used in long-term forecasting papers like [28].
Monthly.  We choose TourismL (Tourism Large) [21] as one of the target datasets. It contains monthly tourist visit data in Australia that has been grouped into various regions. It consists of 555 time-series with very different scales. We choose the task of predicting a 12 month horizon given a context length of 64.
All the target datasets are divided into train:validation:test splits (periods) chronologically with the proportions being 7:1:2. We evaluate the models based on metrics resulting from rolling windows in the test period. Specifically, PreDcT(ZS) solely predicts in the test period as it has already been pretrained. The supervised learning models (per dataset) are trained on the train part with the hyper-parameters being tuned using the validation split. Then they predict in the test period for a head-to-head comparison with the zero-shot PreDcT(ZS).
Baselines. We compare our model against three recently published state-of-the-art supervised forecasting models PatchTST [13], TiDE [5] and FEDFormer [25] as well as the popular N-BEATS model [14]. Note that these models have already been shown [25, 28] to be superior to common statistical forecasting methods such as Prophet [19] and ARIMA; hence we do not include them in our baselines. We train the above supervised models on the train split of each target dataset and measure their performance on the corresponding test split. For these supervised models we report the best metrics among models trained with and without date-derived features. See Appendix A.2 for the hyper-parameters used for each dataset. We compare our zero-shot metrics from PreDcT(ZS) to these state-of-the-art supervised metrics, which we denote by PatchTST(S), TiDE(S), N-BEATS(S) and FEDFormer(S) in our results below.
In PreDcT(ZS), we set input_patch_len=32, output_patch_len=128. We train a model with about 225M parameters that uses 16 head multi-head attention in each transformer layer.
Results. In Table 2 we present the main results on all our target datasets. We report normalized metrics NRMSE and WAPE, that are proportional to MSE and MAE, and are defined (for each time series) as
$\mathrm{NRMSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\sqrt{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}},$
$\mathrm{WAPE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{1}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}}.$
The metrics in the table are across all time-series in the test period of the target datasets. The metrics are calculated over all rolling window (context, horizon) pairs that can be extracted from the test period. Note that FEDformer training was not able to scale to the Wiki dataset (the largest of our target datasets) and hence its metrics for Wiki in the table are left blank.
For each column in the table (corresponding to a metric computed for a dataset), we color-code the best performance by blue, the second-best performance by green, and the worst performance by red. We observe that PreDcT(ZS) obtained uniformly good results (in the ballpark of the best supervised models) for a majority of the target datasets. In particular, PreDcT(ZS) obtained the best performance for the TourismL dataset, and close to the best performing models for Wiki, ETTh1 and ETTh2. It was among the best or second-best performing model for $8$ out of the $14$ columns in the table, and the worst-performing model for only one of the columns(NRMSE on ILI). This is particularly remarkable since we use a single, pretrained model evaluated in zero-shot manner on the target datasets, and are comparing here against state-of-the-art supervised baselines trained separately on each of the datasets.
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 0.380 0.245 0.161 0.571 0.307 0.180 0.114 0.115 0.081 0.414 0.132 0.590 0.203 TiDE(S) 0.663 0.374 0.245 0.161 0.588 0.320 0.186 0.120 0.103 0.070 0.455 0.131 0.574 0.194 N-BEATS(S) 0.687 0.421 0.235 0.152 0.608 0.320 0.179 0.111 0.103 0.069 0.410 0.137 0.666 0.213 FEDFormer(S) 0.675 0.411 0.294 0.205 0.635 0.381 0.240 0.152 - - 0.441 0.164 1.080 0.266 PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192
Table 2: We present NRMSE and WAPE metrics for (S)upervised models and our (Z)ero-(S)hot model. The supervised models are trained, tuned and evaluated on the specific target datasets. PreDcT(ZS) metrics are reported on zero-shot performance i.e the model has never seen the target dataset prior to inference. The best number in each column is colored blue, and the second-best number is colored green. The worst performing model metric per column is colored red. It can be seen that the PreDcT(ZS) metrics are uniformly good across all datasets and was among the best or second-best performing model for $8$ out of the $14$ columns. We report standard errors for the supervised metrics in Table 7 in the appendix.
5.2 Ablation
Next, we perform several ablation studies that inform the design decisions we made for our foundation model architecture.
Different architectures on same pretraining data. [13] have shown that PatchTST can be used to learn semi-supervised representations of time-series. Similarly  [15] have shown that the N-BEATS architecture can be used for transfer learning in time-series. Therefore, we also consider pre-training a foundation model based on the PatchTST and N-BEATS architecture using the same pretraining dataset, and evaluate them in a zero-shot manner on the target datasets, similar to PreDcT(ZS). These baselines will be denoted by PatchTST(ZS) and N-BEATS(ZS). Note that N-BEATS(ZS) was restricted to training and inference with a fixed context length on account of being a MLP model.
The results are shown in Table 3. It can be seen that PreDcT(ZS) performs better than or similar to PatchTST(ZS) on ETTh1, ETTh2, ETTm2, Wiki and TourismL, but we are dramatically better than PatchTST(ZS) on ETTm2 and ILI. Note that because of encoder-decoder only training PatchTST can only adapt to context lengths that are used for pretraining which are 512, 256 and 64 as mentioned in the Pretraining Data section above. This is evident by its bad performance on the ILI dataset which has a context length of 96. This can be further seen in the study in Table 4, discussed subsequently. N-BEATS(ZS) performs slightly better than us on ETTm2, slightly worse on ETTm1, and similar on ETTh1 and ETTh2. But it cannot adapt to varying context lengths, so it could not generalize to the other datasets for Wiki, ILI and TourismL.
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192 PatchTST(ZS) 0.670 0.374 0.239 0.151 0.740 0.370 0.205 0.127 0.109 0.073 0.618 0.199 0.505 0.186 N-BEATS(ZS) 0.670 0.381 0.239 0.153 0.617 0.323 0.189 0.119 - - - - - -
Table 3: We present metrics for the three different zero-shot model architectures. It can be see that PreDcT(ZS) does uniformly well across all datasets. PatchTST(ZS) does not do well on ILI on account of not being able to generalize to context 96 because of encoder-decoder mode of training. N-BEATS numbers could not be obtained on non-ETT datasets because it has a fixed context length due to its MLP architecture. The best number in each column is made bold.
Adapting to different context lengths.  A good foundation model should be able to adapt to a variety of different context lengths. This is possible in our model because of decoder-only training - the output token of every patch extracts features from all the patches that come before it, and is trained to predict the next output patch. In Table 4 we show the performance (in terms of NRMSE) of PreDcT(ZS) with different context lengths on the same task as before of predicting 96 time-points. We also juxtapose our performance with PatchTST(ZS) that is trained in encoder-decoder fashion. It can be seen that our model has good performance throughout which becomes progressively better with more context. On the other hand the performance of PatchTST is only good for context length 512 because it has not been optimized for other context lengths because of encoder-decoder model training. Note that because of overlapping stride the original PatchTST model does not lend itself easily to decoder-only training.
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 96 192 384 512 96 192 384 512 96 192 384 512 96 192 384 512 PreDcT(ZS) 0.779 0.715 0.692 0.672 0.253 0.250 0.239 0.238 0.663 0.631 0.607 0.591 0.204 0.202 0.201 0.199 PatchTST(ZS) 1.429 1.125 0.995 0.670 0.333 0.279 0.245 0.239 1.168 1.327 1.106 0.740 0.320 0.261 0.220 0.205
Table 4: NRMSE numbers are presented for the pretrained models when the context length is varied at inference time. The prediction horizon is held fixed at 96. It can be seen that PreDcT(ZS) can adapt to different context lengths at inference time.
Input patch length.  The size of input_patch_len represents an important trade-off. We have typically seen that increasing its value from 8 to 64 increases performance but having too high a input_patch_len is impractical because the model cannot be easily applied to context lengths that are less than input_patch_len, at inference time. In many monthly and higher granularity tasks, it is common to have small context lengths. In Table 5 we show the NRMSE of another PreDcT(ZS) model with input_patch_len=8 on ETT datasets, which is clearly worse than our original model that uses input_patch_len=32.
input_patch_len ETTh1 ETTh2 ETTm1 ETTm2
32 0.672 0.238 0.591 0.199
8 0.680 0.263 0.706 0.209
Table 5: Ablation with respect to input patch length. NRMSE numbers are reported.
Autoregressive decoding. In recent long-term forecasting works [24, 13, 5] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks. For a foundation model the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons. However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps. This was one of the key decisions in the design of PreDcT, that is quite different from LLMs. In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets. In Table 6, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128. The former has to perform 16 autoregressive steps while the latter has to do only 4. It can be clearly seen that having a larger output_patch_len helps in this case.
output_patch_len ETTh1 ETTh2 ETTm1 ETTm2
128 0.724 0.286 0.711 0.256
32 0.746 0.292 0.737 0.259
Table 6: Ablation with respect to output patch length for the task of predicting 512 steps into the future. NRMSE numbers are reported.
6 Discussion and Future Work
We train a decoder-only foundation model for time-series forecasting using a large pretraining corpus of about 1B time-points, the majority of it being search interest time-series derived from Google trends. We show that even a relatively small 225M parameter pretrained model that uses our PreDcT architecture displays impressive zero-shot performance on a variety of public benchmarks from different domains and granularities. The PreDcT(ZS) model can rival the performance of recent state-of-the-art supervised baselines that have been specifically trained on the target datasets. This is remarkable since the PreDcT(ZS) model has not seen the target datasets before inference.
In future work, it would be interesting to push the boundaries of scale in both the pretraining data as well as the number of parameters in the model. It would be a insightful to perform a scaling study in the lines of [7] for time-series foundation model. An empirical analysis of probabilistic zero-shot forecasting from an appropriately trained foundation model is also an interesting direction for future work.
Appendix A Appendix
A.1 Date Derived Features
In our study we focus on 5 date derived features: (1) month of the year, (2) day of the week, (3) hour of the day, (4) minute of the hour and (5) second of the minute. For any time point $t$ in the context and the horizon we have $\mathbf{x}_{t}\in\mathbb{R}^{5}$. In addition for each date derived feature:
  1. 1.
    We mask it by -1 if its granularity is irrelevant to the time series $\mathbf{y}$. For example, the daily Wikipedia makes no use of hour of the day, minute of the hour, and second of the minute.
  2. 2.
    We normalize its values to $[-0.5,0.5]$. For example, the raw minute of the hour value $v$ is transformed to $v/60-0.5$.
A.2 More Details on Models
N-BEATS. All N-BEATS(S) and N-BEATS(ZS) models are trained following the N-BEATS-G hyper-parameters used in Table-18 from the original paper [14], except that we use width = 1024 and stacks = 24 to match the model sizes to about 200M. All supervised N-BEATS(S) use the dataset specific input output lengths, while N-BEATS(ZS) is trained with input length = 512 and output length = 96 to predict on all ETT datasets.
PatchTST. For PatchTST(S) models in all ETT datasets and ILI we use the hyper-parameters from the original paper [13]. For Wiki and TourismL, we use 16 attention heads, 10 layers, a patch-size of 64 and a stride of 32. The model dimension used is 1024. We tune the learning rate per dataset.
For PatchTST(ZS) we use we use 16 attention heads, 20 layers, a patch-size of 32 and a stride of 16. We use the same model dimension as PreDcT(ZS).
TiDE. We used the hyper-parameters from the original paper [5], but tuned the model dimensions (from 256 to 1024), number of layers (from 1 to 4) and learning date per-dataset.
FEDFormer. For all datasets we use the same hyper-parameters of the original paper [25], except model dimensions. We tuned the model dimension per dataset from 64 to 1024.
PreDcT. We use use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280.
A.3 Additional Empirical Results
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 $\pm$ 0.0004 0.380 $\pm$ 0.0002 0.245 $\pm$ 0.0004 0.161 $\pm$ 0.0004 0.571 $\pm$ 0.003 0.307 $\pm$ 0.001 0.180 $\pm$ 0.0008 0.114 $\pm$ 0.0007 TiDE(S) 0.663 $\pm$ 0.0001 0.374 $\pm$ 0.0001 0.245 $\pm$ 0.002 0.161 $\pm$ 0.001 0.588 $\pm$ 0.0005 0.320 $\pm$ 0.0005 0.186 $\pm$ 0.0005 0.120 $\pm$ 0.0001 N-BEATS(S) 0.687 $\pm$ 0.002 0.421 $\pm$ 0.002 0.235 $\pm$ 0.0002 0.152 $\pm$ 0.0001 0.608 $\pm$ 0.007 0.320 $\pm$ 0.004 0.179 $\pm$ 0.001 0.111 $\pm$ 0.0003 FEDFormer(S) 0.675 $\pm$ 0.003 0.411 $\pm$ 0.004 0.294 $\pm$ 0.010 0.205 $\pm$ 0.01 0.635 $\pm$ 0.007 0.381 $\pm$ 0.005 0.240 $\pm$ 0.006 0.152 $\pm$ 0.004 Model Wiki ILI TourismL - NRMSE WAPE NRMSE WAPE NRMSE WAPE - - PatchTST(S) 0.115 $\pm$ 0.0002 0.081 $\pm$ 0.0002 0.414 $\pm$ 0.005 0.132 $\pm$ 0.003 0.595 $\pm$ 0.012 0.204 $\pm$ 0.003 - - TiDE(S) 0.103 $\pm$ 0.0004 0.070 $\pm$ 0.0006 0.455 $\pm$ 0.004 0.131 $\pm$ 0.005 0.574 $\pm$ 0.001 0.194 $\pm$ 0.002 - - N-BEATS(S) 0.103 $\pm$ 0.0002 0.069 $\pm$ 0.0001 0.410 $\pm$ 0.003 0.137 $\pm$ 0.002 0.666 $\pm$ 0.002 0.213 $\pm$ 0.0003 - - FEDFormer(S) - - 0.441 $\pm$ 0.0007 0.164 $\pm$ 0.0003 1.080 $\pm$ 2 0.266 $\pm$ 0.0007 - -
Table 7: NRMSE and WAPE confidence intervals (mean $\pm$ 1 standard error) of the supervised baselines.

### v2

@@
 Abstract
+Author names are listed in alphabetical order.1. Google Research, Correspond to senrajat@google.com
 Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset.
-Our model is based on pretraining a patched-decoder style attention model on a large timeseries corpus, and can work well across different forecasting history lengths, prediction lengths and temporal granularities.
-Author names are listed in alphabetical order.
+Our model is based on pretraining a patched-decoder style attention model on a large time-series corpus comprising both real-world and synthetic datasets.
+Experiments on a diverse set of previously unseen forecasting datasets suggests that the model can yield accurate zero-shot forecasts across different domains, forecasting horizons and temporal granularities.
+Machine Learning, ICML
+[…]
+[…]
 1 Introduction
@@
 In spite of these issues, we provide evidence to answer the above question in the affirmative.
-In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets with different temporal granularities, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets).
+In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets).
 Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time.
-The key elements of our foundation model are twofold: 1) a time series corpus built mostly using Google Trends [^1] https://trends.google.com, that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus.
-Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (225M parameters) and pretraining data size (1B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data.
+The key elements of our foundation model are twofold: 1) a large-scale time series corpus built using both synthetic and real-world (mostly using time series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/), that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus.
+Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (100B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data.
+Our work also suggests that unlike recent work [9] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a fraction of its costs.
 2 Related Work
@@
 For a more in-depth discussion about transfer learning in time-series we refer the reader to the survey in [16].
-[26] show how to use the the GPT-2 backbone [16] for various tasks including time-series forecasting.
-[4] is a follow up works along the same lines.
+There has been some recent work on re-using or fine-tuning large language models for time-series forecasting.
+In particular,  [9] benchmarks pretrained LLMs like GPT-3 and LLaMA-2 for zero-shot forecasting performance.
+As we show later, our model obtains much superior zero-shot performance at a tiny fraction of these model sizes.
+Zhou et al.
+[36] and [5] show how to fine-tune a GPT-2 [24] backbone model for time-series forecasting tasks.
 Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset.
-For instance Table-18 [26] shows M4 to M3 transfer.
-The rest of the two papers are mostly focused on fine-tuning and to the best of our knowledge they do not train a single foundation model that shows out of the box zero-shot performance on a variety of datasets.
-To the best of our knowledge, the very recent work in TimeGPT-1 [6] is the only known parallel work on foundation model for time-series.
+With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on a target dataset, and not on pretraining a single foundation model that shows out-of-the box zero-shot performance on a variety of datasets.
+To the best of our knowledge, the very recent work in TimeGPT-1 [7] is the only other parallel work on a zero-shot foundation model for time-series forecasting.
 However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
@@
 The task at hand is to build a general purpose zero-shot forecaster that takes in the past $C$ time-points of a time-series as context and predicts the future $H$ time-points.
-Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy like notation for indices.
+Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy-like notation for indices.
 Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$.
-Note that since we are building a one-fits-all model we cannot have dataset specific dynamic or static covariates during training time.
+Note that since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time.
 However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$.
-See Appendix A.1 for details.
 Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$.
@@
 For instance, if the metric is Mean Squared Error (MSE), then the goodness of fit is measured by,
-$\mathrm{MSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)=\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}.$ (2)
+$\mathrm{MSE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)$
+$=\frac{1}{H}\mathopen{}\mathclose{{\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}}\right\|_{2}^{2}.$ (2)
 4 Model Architecture
@@
 Transformers have been shown to be able to adapt to different context lengths in NLP [24].
-Inspired by the success of patch based modeling in the recent long horizon forecasting work [13] we also chose to breakdown the time-series into patches during training.
-However, there are several key differences in our foundation model architecture, the primary one being that our model is trained in decoder-only mode [9].
-We will now describe the key parts of our architecture and training methodology illustrated in Figure 1.
-[…]
-[width=]fts_new_v2.pdf
+However, there are several time-series specific design choices.
+The main guiding principles for our architecture are the following:
+Patching.
+Inspired by the success of patch based modeling in the recent long horizon forecasting work [19] we also choose to breakdown the time-series into patches during training.
+A patch of a time-series is a natural analogues for a token in language models and have been shown to improve performance.
+Moreover this improves inference speed as the number of tokens being fed into the transformer is reduced by a factor of the patch length.
+On the other hand, increasing the patch length all the way to the context length moves us away from decoder-only training and the efficiencies that come with it.
+We delve into this further in Section 6.2.
+Decoder-only model.
+A key difference between our architecture and PatchTST [19] is that our model is trained in decoder-only mode [15].
+In other words, given a sequence of input patches the model is optimized to predict the next patch as a function of all past patches.
+Similar to LLM’s this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
+Longer output patches.
+In LLM’s the output is always generated in an auto-regressive fashion one token at a time.
+However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy that multi-step auto-regressive decoding [33].
+But the latter is not possible when the horizon length is not known a prior like in the case of zero-shot forecasting, which is our primary goal.
+We propose a middle ground by allowing the output patches for prediction to be be longer than the input patches.
+As an example, suppose the input patch length is 32 and output patch length is 128.
+During training, the model is simultaneously trained to use the first 32 time-points to forecast the next 128 time-steps, the first 64 time-points to forecast time-steps 65 to 192, the first 96 time-points to forecast time-steps 97 to 224 and so on.
+During inference, suppose the model is given a new time-series of length 256 and tasked with forecasting the next 256 time-steps into the future.
+The model will first generate the future predictions for time-steps 257 to 384, then condition on the initial 256 length input plus the generated output to generate time-steps 385 to 512.
+On the other hand, if in our model the output patch length was equal to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above.
+Patch Masking.
+If we use patches naively, the model might only learn to predict well for context lengths that are multiples of the input patch length.
+Therefore we make a careful use of masking during training.
+Parts of patches as well as entire patches from the beginning of the context window can be masked in a data batch.
+We employ a specific random masking strategy during training that helps to model see any all possible context lengths starting from 1 to a maximum context length.
+Now that we have mentioned the guiding principles, we next formally describe each component of our architecture that is illustrated in Figure 1.
 Figure 1: We provide an illustration of our model architecture during training where we show a input time-series of a certain length that can be broken down into input patches.
@@
 We first break the input into contiguous non-overlapping patches.
-Then each patch (along with optional date derived features for that patch) is processed by a Residual Block into a vector of size model_dim.
+Then each patch is processed by a Residual Block into a vector of size model_dim.
+Along with the input, we also supply a binary padding mask $\mathbf{m}_{1:L}$ where 111 denotes that the corresponding input in $\mathbf{y}_{1:L}$ should be ignored and vice-versa.
 The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [6].
-In other words, the inputs $\mathbf{y}_{1:L},{\bm{x}}_{1:L}$ are broken down into patches of size input_patch_len ($p$).
-The $j$-th patch can be denoted as $\tilde{\mathbf{y}}_{j}=\mathbf{y}_{p(j-1)+1:pj}$ and $\tilde{\mathbf{x}}_{j}=\mathbf{x}_{p(j-1)+1:pj}$.
+In other words, the inputs $\mathbf{y}_{1:L}$ are broken down into patches of size input_patch_len ($p$).
+The $j$-th patch can be denoted as $\tilde{\mathbf{y}}_{j}=\mathbf{y}_{p(j-1)+1:pj}$.
+Similarly the mask can also be patched as $\tilde{\mathbf{m}}_{j}=\mathbf{m}_{p(j-1)+1:pj}$.
 Then the $j$-th input token to the subsequent transformer layers can be denoted as,
-$\mathbf{t}_{j}=\texttt{InputResidualBlock}(\tilde{\mathbf{y}}_{j},\tilde{\mathbf{x}}_{j})+\texttt{PE}_{j}$ (3)
+$\mathbf{t}_{j}=\texttt{InputResidualBlock}(\tilde{\mathbf{y}}_{j}\odot(1-\tilde{\mathbf{m}}_{j}))+\texttt{PE}_{j}$ (3)
 where $\texttt{PE}_{j}$ denotes the $j$-th positional encoding as defined in the original transformer paper [28].
@@
 Stacked Transformer.
-The bulk of the parameters in our model are in $n_{l}$ transformer layers stacked on top of each other.
-Each of these layers have the standard multi-head self-attention (MHA) followed by a feed-forward network (FFN).
+The bulk of the parameters in our model are in num_layers ($n_{l}$) transformer layers stacked on top of each other.
+Each of these layers have the standard multi-head self-attention (SA) followed by a feed-forward network (FFN).
 The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads).
@@
 This can be described by the equation
-$\mathbf{o}_{j}=\texttt{StackedTransformer}\mathopen{}\mathclose{{}\left(\mathbf{t}_{1},\cdots,\mathbf{t}_{j}}\right),~{}~{}~{}\forall j\in[N].$ (4)
+$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$
+$\forall j\in[N].$ (4)
+In Equation (4), $\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked.
+All patches that are masked out completely are not attended to by the causal self attention.
 Output Layers.
@@
 Another approach can be to output the logits of a probability distribution family and minimize the maximum likelihood loss for probabilistic forecasting [1, 25].
+Training.
+We train the model with standard mini-batch gradient descent in decoder only fashion, that goes through all windows for a time-series and across time-series.
+The only non-standard part is the way we sample the mask during training.
+For each time-series in the batch, we sample a random number $r$ between 0 and $p-1$.
+Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask our a fraction of the first input patch.
+However, this is sufficient to cover all input context lengths from 1 to the maximum training context length.
+Let us say that the maximum context length is 512 and $p=32$.
+Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized predict after seeing 28=32428=32-428 = 32 - 4 time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing 28+3228+3228 + 32 time-points, and so on.
+When this argument is repeated for all such $r$’s, we can see that the model has seen all possible context lengths from 111 to 512512512.
 Inference.
@@
 Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on.
-We name our model Pretrained Decoder for Time-series (PreDcT).
-5 Empirical Results
-We evaluate our model in zero-shot settings on well known public datasets against state-of-the-art supervised forecasting baselines.
-We show that a single pretrained model can come close or surpass the performance of baselines on the benchmarks even when the baselines are specially trained or tuned for each specific task.
-Subsequently, we perform ablation studies that justify different choices made in our foundation model architecture.
-5.1 Zero-shot Evaluation
-We are interested evaluating the performance of a single foundation model pretrained on a large time-series dataset on different target datasets never seen before by the model or, in other words, zero-shot evaluation.
-To test the generalization performance of our baselines, we choose target datasets covering a variety of data domains, scales and time granularities.
-We will first discuss the pretraining and target datasets, along with the baselines we use.
-Then we demonstrate the (Z)ero-(S)hot performance of our model PreDcT(ZS) compared against state-of-the-art models trained directly on the target datasets.
-Pretraining Data.
+If $L$ is not a multiple of $p$ we can append some zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.
+We name our Time-series Foundation Model as TimesFM.
+5 Pretraining Details
 We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model.
 It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model.
-In this paper, we find that Google Trends [^2] https://trends.google.com can provide a time series corpus that is ideally suited for pre-training our foundation model.
+The bulk of real data used to train our models come from Google trends  [^3] https://trends.google.com and Wiki Pageview statistics  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics.
+In summary the main data sources are:
+Google Trends.
 Google Trends captures search interest over time for millions of queries.
@@
 The date ranges are Jan.
-2018 to Dec 2019 for hourly and Jan.
+2018 to Dec.
+2019 for hourly and Jan.
 2007 to Dec.
 2021 for the other granularities.
-Along with the trends data, we also add time series from several other publicly available datasets to our pretraining corpus.
+Wiki Pageviews.
+Wiki Pageviews captures the hourly views of all Wikimedia pages.
+We download all pageview data from Jan.
+2012 to Nov.
+2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out the resulting time-series with excessive zeros.
+The final corpus contains O(100B) time points.
+Other real data sources.
+Along with the wiki and trends data, we also add time-series from several other publicly available datasets to our pretraining corpus.
 We add in all the granularities of the M4 dataset [17] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [34]).
+We also add the 10 minute granularity Weather dataset used for evaluations in [34].
 M4 has a good mix of granularities with around 100k time-series in total.
@@
 In addition, we add all the 15 min granularity traffic time series from [29].
-We train on a mixture distribution over these datasets that aims to give sufficient weightage to all granularities.
+Synthetic.
+Another major component of our pretraining data is of synthetic origin.
+We create generators for ARMA [18] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions.
+A synthetic time-series can be an additive combination of 1 or more of these processes.
+We create 3M synthetic time-series each of length 2048 time-points.
+More details are presented in Appendix A.4.
+Dataset Mixing.
+We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities.
+The training mixture is 50% real data and 50% synthetic, with the real data mixture providing equal weights to all hourly + sub-hourly, daily, weekly, and monthly datasets.
 We train with a maximum context length of 512 whenever the length of the time-series allows that.
-For weekly granularity we do not have sufficiently long time-series therefore a max.
-context length of 256 is used.
-For the same reason, max.
-context length of 64 is used while training on monthly and higher granularity data.
-Target Datasets.
-To benchmark our model’s performance, we choose commonly used forecasting datasets of varying sizes that cover various domains, granularities, context lengths and horizon lengths, to test the generalization power of our foundation model against other baselines.
-The details are summarized in Table 1.
-Dataset # Time-Series Time Length Granularity Context Horizon
-ETTm1 7 69680 15 Min. 512 96
-ETTm2 7 69680 15 Min. 512 96
-ETTh1 7 17420 1 Hour 512 96
-ETTh2 7 17420 1 Hour 512 96
-Wiki 115084 803 1 Day 256 56
-ILI 7 966 1 Week 96 24
-TourismL 555 228 1 Month 64 12
-Table 1: Statistics and task definitions of the target datasets.
-(Sub)Hourly.
-For 15 min.
-granularity we use the ETTm1, ETTm2 datasets and test all models on the task of predicting a horizon of 96 time-points after seeing a context of size 512.
-For hourly granularity, we choose the ETTh1, ETTh2 datasets and test all models on the same task as the 15 min.
-datasets.
-The datasets and the context, horizon pair have been widely used in long-term forecasting benchmarks [28, 13].
-Note that we used the more challenging original, unscaled versions of these datasets in order to test our model’s zero-shot performance on time-series of different scales.
-Daily.
-For the daily granularity, we use the Wikipedia web-traffic dataset from the corresponding Kaggle competition  [^3] https://www.kaggle.com/code/muonneutrino/wikipedia-traffic-data-exploration.
-It has 115k time-series with over two years of data if we exclude the time-series with missing values.
-The dataset contains web-traffic to Wikipedia articles and the task is to predict the web-traffic (in $\log$ scale) on future dates.
-We choose a context length of 256 to predict a horizon of 56 days (8 weeks).
-This dataset has been used in prior multivariate forecasting papers such as like [18].
-Weekly.
-We use the ILI dataset [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html that collects the number of patients and influenza-like illness ratio in a weekly frequency.
-We use a context length of 96 to predict 24 weeks into the future.
-This is one of the configurations used in long-term forecasting papers like [28].
-Monthly.
-We choose TourismL (Tourism Large) [21] as one of the target datasets.
-It contains monthly tourist visit data in Australia that has been grouped into various regions.
-It consists of 555 time-series with very different scales.
-We choose the task of predicting a 12 month horizon given a context length of 64.
-All the target datasets are divided into train:validation:test splits (periods) chronologically with the proportions being 7:1:2.
-We evaluate the models based on metrics resulting from rolling windows in the test period.
-Specifically, PreDcT(ZS) solely predicts in the test period as it has already been pretrained.
-The supervised learning models (per dataset) are trained on the train part with the hyper-parameters being tuned using the validation split.
-Then they predict in the test period for a head-to-head comparison with the zero-shot PreDcT(ZS).
-Baselines.
-We compare our model against three recently published state-of-the-art supervised forecasting models PatchTST [13], TiDE [5] and FEDFormer [25] as well as the popular N-BEATS model [14].
-Note that these models have already been shown [25, 28] to be superior to common statistical forecasting methods such as Prophet [19] and ARIMA; hence we do not include them in our baselines.
-We train the above supervised models on the train split of each target dataset and measure their performance on the corresponding test split.
-For these supervised models we report the best metrics among models trained with and without date-derived features.
-See Appendix A.2 for the hyper-parameters used for each dataset.
-We compare our zero-shot metrics from PreDcT(ZS) to these state-of-the-art supervised metrics, which we denote by PatchTST(S), TiDE(S), N-BEATS(S) and FEDFormer(S) in our results below.
-In PreDcT(ZS), we set input_patch_len=32, output_patch_len=128. We train a model with about 225M parameters that uses 16 head multi-head attention in each transformer layer.
-Results. In Table 2 we present the main results on all our target datasets. We report normalized metrics NRMSE and WAPE, that are proportional to MSE and MAE, and are defined (for each time series) as
-$\mathrm{NRMSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\sqrt{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}},$
-$\mathrm{WAPE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{1}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}}.$
-The metrics in the table are across all time-series in the test period of the target datasets.
-The metrics are calculated over all rolling window (context, horizon) pairs that can be extracted from the test period.
-Note that FEDformer training was not able to scale to the Wiki dataset (the largest of our target datasets) and hence its metrics for Wiki in the table are left blank.
-For each column in the table (corresponding to a metric computed for a dataset), we color-code the best performance by blue, the second-best performance by green, and the worst performance by red.
-We observe that PreDcT(ZS) obtained uniformly good results (in the ballpark of the best supervised models) for a majority of the target datasets.
-In particular, PreDcT(ZS) obtained the best performance for the TourismL dataset, and close to the best performing models for Wiki, ETTh1 and ETTh2.
-It was among the best or second-best performing model for $8$ out of the $14$ columns in the table, and the worst-performing model for only one of the columns(NRMSE on ILI).
-This is particularly remarkable since we use a single, pretrained model evaluated in zero-shot manner on the target datasets, and are comparing here against state-of-the-art supervised baselines trained separately on each of the datasets.
-[…]
-!
-Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 0.380 0.245 0.161 0.571 0.307 0.180 0.114 0.115 0.081 0.414 0.132 0.590 0.203 TiDE(S) 0.663 0.374 0.245 0.161 0.588 0.320 0.186 0.120 0.103 0.070 0.455 0.131 0.574 0.194 N-BEATS(S) 0.687 0.421 0.235 0.152 0.608 0.320 0.179 0.111 0.103 0.069 0.410 0.137 0.666 0.213 FEDFormer(S) 0.675 0.411 0.294 0.205 0.635 0.381 0.240 0.152 - - 0.441 0.164 1.080 0.266 PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192
-Table 2: We present NRMSE and WAPE metrics for (S)upervised models and our (Z)ero-(S)hot model.
-The supervised models are trained, tuned and evaluated on the specific target datasets.
-PreDcT(ZS) metrics are reported on zero-shot performance i.e the model has never seen the target dataset prior to inference.
-The best number in each column is colored blue, and the second-best number is colored green.
-The worst performing model metric per column is colored red.
-It can be seen that the PreDcT(ZS) metrics are uniformly good across all datasets and was among the best or second-best performing model for $8$ out of the $14$ columns.
-We report standard errors for the supervised metrics in Table 7 in the appendix.
+For weekly granularity we do not have sufficiently long time-series therefore a maximum context length of 256 is used.
+For the same reason, a maximum context length of 64 is used while training on $\geq$ monthly granularity data.
+(a) Monash Archive [8]
+(b) Darts [10]
+(c) ETT (Horizons 96 and 192) [34]
+Figure 2: We report average performance in three groups of datasets.
+In all figures, the lower the metric the better and the error bars represent one standard error.
+Note that among the baselines onlye TimesFM and llmtime are zero-shot.
+In (a) we report results on the Monash datasets.
+Since the datasets have different scales, we average the MAE’s scaled by the MAE of a naive baseline.
+We can see that TimesFM is among the top-3 models.
+In (b), we report the similarly scaled MAE on the Darts benchmarks.
+TimesFM is within significance of the top of the best performing method which is ARIMA in this case.
+Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones.
+Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total.
+TimesFM and PatchTST are the best performing models in this case.
+6 Empirical Results
+We evaluate our model in zero-shot settings on three groups of well known public datasets against the best performing baselines for each group.
+These datasets have been intentionally held out from our pretraining data.
+We show that a single pretrained model can come close or surpass the performance of baselines models on the benchmarks even when the baselines are specially trained or tuned for each specific task.
+Subsequently, we perform ablation studies that justify different choices made in our architecture.
+6.1 Zero-shot Evaluation
+To benchmark our model’s performance, we choose three groups of commonly used forecasting datasets that cover various domains, sizes, granularities, and horizon lengths: Darts [10], Monash [8] and Informer datasets [34], to test the generalization power of our foundation model against other baselines.
+In all cases, we report performance on the official metrics and scalings of the datasets. We present a summary of the results below - more detailed experimental results can be found in Appendix A.2.
+Monash [8].
+Monash archive is a collection of 30 datasets of different training and prediction lengths that covers granularities ranging from minutes to years and domains including finance, demand forecasting, weather and traffic.
+The archive reports four official metrics for several statistical baselines such as Exponential Smoothing(ETS) and ARIMA, as well as supervised ML baselines like CatBoost [23], DeepAR [25] and WaveNet [20].
+Following llmtime [9] we start from the Monash Huggingface repository  [^5] https://huggingface.co/datasets/monash_tsf and filter out the datasets that contain missing values.
+This leaves us with 18 datasets.
+Out of the four official metrics, following prior work [9], we report our performance in terms of mean MAE (see Appendix A.1).
+As the datasets have massively different scales, for each dataset we normalize the metric by the metric achieved by a naive baseline that just constantly predicts the last value in the context for each time-series.
+Then the scaled MAE’s are averaged across all datasets.
+The scaled aggregation was also used in [9].
+The mean scaled MAE across all datasets is plotted in Figure 2(a) along with standard error bars.
+We compare the performance of TimesFMwith the baseline models implemented in Monash, and the zero-shot llmtime [9] model that uses GPT-3 [24] with a specific prompting technique.
+Note that the zero-shot models are marked as (Zero-Shot).
+TimesFM is among the top 3 models even though we never trained on these datasets.
+It outperforms some deep supervised models like DeepAR and FFNN, and improves on llmtime’s performance by more than 10%.
+Darts [10].
+This is a collection of 8 univariate datasets which include interesting seasonalities and additive+multiplicative trends.
+We report performance of several baselines implemented in the Darts package like TCN [14], N-HiTS [4] and N-BEATS [21].
+All these baselines are supervised.
+As before, we also report zero-shot forecasting results from llmtime [9] using GPT-3 [24].
+Other supervised baselines implemented in [9] like SM-GP [31] and seasonal ARIMA [18] are also added.
+We report the official metric for this dataset group that is MAE for each individual dataset in Appendix A.2.
+In Figure 2(b), we present the average scaled MAE across all 8 datasets, as we did for the Monash datasets.
+TimesFM is within statistical significance of the best model that is seasonal ARIMA in this case.
+Note that since there are only 8 individual time-series in this dataset group, the standard errors are not sharp and therefore does not provide a clear ordering among the models.
+Informer [34].
+The Informer datasets have been widely used for benchmarking various supervised long-horizon forecasting methods.
+A few of these datasets are used in pretraining, so we focus on the other datasets in this collection (ETTm1, ETTm2.
+ETTh1 and ETTh2) related to electricity transformer temperatures over a two year period in 111 hour and 151515 minutes granularities.
+Note that the long horizon baselines usually report rolling validation results on the test set which would amount to millions of tokens for evaluating llmtime [9] and would be too expensive.
+Therefore, following the latter paper, we compare all methods on the last test window.
+Also, it is reasonable to directly average the MAE for these datasets since the results are reported on standard normalized dataset (using the statistics of the training portion).
+We consider the task of predicting horizon length 96 and 192, given a context length of 512 for all methods.
+The MAE averaged over all 8 tasks (4 datasets with two horizons each) is presented in Figure 2(b).
+TimesFM performs the best and the supervised PatchTST [19] baseline (which is a state-of-the-art long horizon deep forecasting method) is within significance of it.
+The other long horizon methods are quite a bit worse even though they have been trained these datasets.
+llmtime(ZS) is better than FEDFormer but worse than PatchTST in a statistically significant way.
 6.2 Ablation
-Next, we perform several ablation studies that inform the design decisions we made for our foundation model architecture.
-Different architectures on same pretraining data.
-[13] have shown that PatchTST can be used to learn semi-supervised representations of time-series.
-Similarly  [15] have shown that the N-BEATS architecture can be used for transfer learning in time-series.
-Therefore, we also consider pre-training a foundation model based on the PatchTST and N-BEATS architecture using the same pretraining dataset, and evaluate them in a zero-shot manner on the target datasets, similar to PreDcT(ZS).
-These baselines will be denoted by PatchTST(ZS) and N-BEATS(ZS).
-Note that N-BEATS(ZS) was restricted to training and inference with a fixed context length on account of being a MLP model.
-The results are shown in Table 3.
-It can be seen that PreDcT(ZS) performs better than or similar to PatchTST(ZS) on ETTh1, ETTh2, ETTm2, Wiki and TourismL, but we are dramatically better than PatchTST(ZS) on ETTm2 and ILI.
-Note that because of encoder-decoder only training PatchTST can only adapt to context lengths that are used for pretraining which are 512, 256 and 64 as mentioned in the Pretraining Data section above.
-This is evident by its bad performance on the ILI dataset which has a context length of 96.
-This can be further seen in the study in Table 4, discussed subsequently.
-N-BEATS(ZS) performs slightly better than us on ETTm2, slightly worse on ETTm1, and similar on ETTh1 and ETTh2.
-But it cannot adapt to varying context lengths, so it could not generalize to the other datasets for Wiki, ILI and TourismL.
-[…]
-!
-Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192 PatchTST(ZS) 0.670 0.374 0.239 0.151 0.740 0.370 0.205 0.127 0.109 0.073 0.618 0.199 0.505 0.186 N-BEATS(ZS) 0.670 0.381 0.239 0.153 0.617 0.323 0.189 0.119 - - - - - -
-Table 3: We present metrics for the three different zero-shot model architectures.
-It can be see that PreDcT(ZS) does uniformly well across all datasets.
-PatchTST(ZS) does not do well on ILI on account of not being able to generalize to context 96 because of encoder-decoder mode of training.
-N-BEATS numbers could not be obtained on non-ETT datasets because it has a fixed context length due to its MLP architecture.
-The best number in each column is made bold.
-Adapting to different context lengths.
-A good foundation model should be able to adapt to a variety of different context lengths.
-This is possible in our model because of decoder-only training - the output token of every patch extracts features from all the patches that come before it, and is trained to predict the next output patch.
-In Table 4 we show the performance (in terms of NRMSE) of PreDcT(ZS) with different context lengths on the same task as before of predicting 96 time-points.
-We also juxtapose our performance with PatchTST(ZS) that is trained in encoder-decoder fashion.
-It can be seen that our model has good performance throughout which becomes progressively better with more context.
-On the other hand the performance of PatchTST is only good for context length 512 because it has not been optimized for other context lengths because of encoder-decoder model training.
-Note that because of overlapping stride the original PatchTST model does not lend itself easily to decoder-only training.
-[…]
-!
-Model ETTh1 ETTh2 ETTm1 ETTm2 96 192 384 512 96 192 384 512 96 192 384 512 96 192 384 512 PreDcT(ZS) 0.779 0.715 0.692 0.672 0.253 0.250 0.239 0.238 0.663 0.631 0.607 0.591 0.204 0.202 0.201 0.199 PatchTST(ZS) 1.429 1.125 0.995 0.670 0.333 0.279 0.245 0.239 1.168 1.327 1.106 0.740 0.320 0.261 0.220 0.205
-Table 4: NRMSE numbers are presented for the pretrained models when the context length is varied at inference time.
-The prediction horizon is held fixed at 96.
-It can be seen that PreDcT(ZS) can adapt to different context lengths at inference time.
-Input patch length.
-The size of input_patch_len represents an important trade-off.
-We have typically seen that increasing its value from 8 to 64 increases performance but having too high a input_patch_len is impractical because the model cannot be easily applied to context lengths that are less than input_patch_len, at inference time.
-In many monthly and higher granularity tasks, it is common to have small context lengths.
-In Table 5 we show the NRMSE of another PreDcT(ZS) model with input_patch_len=8 on ETT datasets, which is clearly worse than our original model that uses input_patch_len=32.
-input_patch_len ETTh1 ETTh2 ETTm1 ETTm2
-32 0.672 0.238 0.591 0.199
-8 0.680 0.263 0.706 0.209
-Table 5: Ablation with respect to input patch length. NRMSE numbers are reported.
+Next, we perform several ablation studies that inform the design decisions we made for our model architecture.
+Scaling.
+Performance curves with respect to number of parameters in a model have been a keenly studied area in the context of LLMs.
+Kaplan et al.
+[12] established a power law like relationship between the number of parameters in a language model and its downstream performance i.e the more the number of paramaters the better the performance.
+However,  Hoffmann et al.
+[11] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset.
+We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pretraining dataset and till the similar number of iterations.
+We provide the mean scaled MAE results on the Monash datasets in Figure 3.
+It can be clearly seen that the errors decrease monotonically with the number of parameters, thus pointing to a similar observation as [12], but in the context of time-series.
+Figure 3: Average scaled MAE on Monash datasets for three different TimesFM model sizes. We can see that the performance improves with increasing model size.
 Autoregressive decoding.
@@
 However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps.
-This was one of the key decisions in the design of PreDcT, that is quite different from LLMs.
+This was one of the key decisions in the design of TimesFM, that is quite different from LLMs.
 In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets.
-In Table 6, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128.
+In Figure 4, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128.
 The former has to perform 16 autoregressive steps while the latter has to do only 4.
 It can be clearly seen that having a larger output_patch_len helps in this case.
-output_patch_len ETTh1 ETTh2 ETTm1 ETTm2
-128 0.724 0.286 0.711 0.256
-32 0.746 0.292 0.737 0.259
-Table 6: Ablation with respect to output patch length for the task of predicting 512 steps into the future. NRMSE numbers are reported.
-6 Discussion and Future Work
-We train a decoder-only foundation model for time-series forecasting using a large pretraining corpus of about 1B time-points, the majority of it being search interest time-series derived from Google trends.
-We show that even a relatively small 225M parameter pretrained model that uses our PreDcT architecture displays impressive zero-shot performance on a variety of public benchmarks from different domains and granularities.
-The PreDcT(ZS) model can rival the performance of recent state-of-the-art supervised baselines that have been specifically trained on the target datasets.
-This is remarkable since the PreDcT(ZS) model has not seen the target datasets before inference.
-In future work, it would be interesting to push the boundaries of scale in both the pretraining data as well as the number of parameters in the model.
-It would be a insightful to perform a scaling study in the lines of [7] for time-series foundation model.
-An empirical analysis of probabilistic zero-shot forecasting from an appropriately trained foundation model is also an interesting direction for future work.
+Figure 4: Ablation with respect to output patch length for the task of predicting 512 steps into the future, in the same setting as that of Section 6.1.
+We report the average across all ETT 4 datasets.
+Input patch length.
+The size of input_patch_len represents an important trade-off.
+We have typically seen that increasing its value from 8 to 32 increases performance but having too high a input_patch_len is impractical since that makes the model shift from decoder only training more towards encoder-decoder style training.
+Note that in the ”Training” paragraph of Section 4, we describe the mask sampling strategy to support any context length.
+If in the extreme case $p$ is set the maximum context length we have to individually sample all possible context windows from 1 to maximum context length, which would be required for encoder-decoder style of training.
+In Figure 5 we show the mean scaled MAE TimesFM(ZS) - 70M model with input_patch_len=8 on Monash datasets, which is clearly worse than our original model that uses input_patch_len=32.
+Figure 5: Average scaled MAE for our 70M models on Monash datasets for two different input patch lengths. We can see that the performance improves with increasing patch size.
+6.3 Visualizing Forecast
+We conduct a visual inspection of the forecasts generated by TimesFM(ZS) on the benchmark datasets.
+Our inspection in particular emphasizes time-series exhibiting both trend and seasonal components readily interpretable by humans.
+As illustrated in Figure 6, TimesFM(ZS) effectively captures these subtle characteristics within both the trend and seasonal patterns of the depicted time-series.
+Figure 6: Forecasts visualized on Darts and Monash datasets. Plots on the right zoom in the forecasting parts from the left.
+7 Conclusion
+In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time series data.
+This model is trained on real-world and synthetic datasets comprising around 100B timepoints using a patched-decoder style attention architecture with around 200M parameters.
+In future work, we plan to delve into a more theoretical understanding of how such a time-series foundation model can obtain good performance for out-of-distribution data, and also investigate the fine-tuning/few-shot performance of this model.
 Appendix A Appendix
-A.1 Date Derived Features
-In our study we focus on 5 date derived features: (1) month of the year, (2) day of the week, (3) hour of the day, (4) minute of the hour and (5) second of the minute.
-For any time point $t$ in the context and the horizon we have $\mathbf{x}_{t}\in\mathbb{R}^{5}$.
-In addition for each date derived feature:
-  1. 1.
-    We mask it by -1 if its granularity is irrelevant to the time series $\mathbf{y}$. For example, the daily Wikipedia makes no use of hour of the day, minute of the hour, and second of the minute.
-  2. 2.
-    We normalize its values to $[-0.5,0.5]$. For example, the raw minute of the hour value $v$ is transformed to $v/60-0.5$.
-A.2 More Details on Models
-N-BEATS.
-All N-BEATS(S) and N-BEATS(ZS) models are trained following the N-BEATS-G hyper-parameters used in Table-18 from the original paper [14], except that we use width = 1024 and stacks = 24 to match the model sizes to about 200M.
-All supervised N-BEATS(S) use the dataset specific input output lengths, while N-BEATS(ZS) is trained with input length = 512 and output length = 96 to predict on all ETT datasets.
-PatchTST.
-For PatchTST(S) models in all ETT datasets and ILI we use the hyper-parameters from the original paper [13].
-For Wiki and TourismL, we use 16 attention heads, 10 layers, a patch-size of 64 and a stride of 32.
-The model dimension used is 1024.
-We tune the learning rate per dataset.
-For PatchTST(ZS) we use we use 16 attention heads, 20 layers, a patch-size of 32 and a stride of 16. We use the same model dimension as PreDcT(ZS).
-TiDE. We used the hyper-parameters from the original paper [5], but tuned the model dimensions (from 256 to 1024), number of layers (from 1 to 4) and learning date per-dataset.
-FEDFormer. For all datasets we use the same hyper-parameters of the original paper [25], except model dimensions. We tuned the model dimension per dataset from 64 to 1024.
-PreDcT. We use use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280.
-A.3 Additional Empirical Results
+A.1 Metrics
+The metrics that are used for reporting results in this paper are:
+  • MAE [8]
+    $\mathrm{MAE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)\\ =\frac{1}{H}\mathopen{}\mathclose{{\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}}\right\|_{1}.$ (7)
+  • msMAPE [8]
+    $\mathrm{msMAPE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)\\ =\frac{1}{H}\sum_{i=1}^{H}\frac{2\lvert y_{L+i}-\hat{y}_{L+i}\rvert}{\max\mathopen{}\mathclose{{\left\{\lvert y_{L+i}\rvert+\lvert\hat{y}_{L+i}\rvert+\epsilon,0.5+\epsilon}}\right\}}.$ (8)
+    In Monash benchmarks [8] $\epsilon=0.1$ was used.
+This metric is used in order to avoid undefined values in other normalized metrics like MAPE.
+In multivariate datasets the metrics are calculated for each time-series and then we take the mean or the median.
+In this paper we only use the mean versions.
+Aggregating across datasets.
+Since the datasets have wildly different scales averaging unnormalized metrics like MAE is not kosher.
+Therefore following [9] we scale the metric of each baseline for a dataset by the same metric achieved by a naive baseline on that dataset.
+The naive baseline just makes the constant prediction $y_{L}$ repeated across the prediction length.
+We did not need to do that for the Informer datasets since on these datasets metrics are usually reported on standard normalized data [19].
+A.2 Additional Empirical Results
+In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1
+A.2.1 Darts
+We present the MAE results individually from all 8 datasets in Table 1.
+It can be seen that TimesFM performs well for all datasets with clear seasonal patterns.
+We do not perform the among the top-3 only in Sunspot and Wooly datasets.
+On an average we are within significant level of the best model and a close second on the mean scaled MAE metric.
+Table 1: MAE for Darts datasets. We also include the naive baseline that predicts the last values in the context repeatedly.
+gp arima TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) NAIVE
+AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 14.75 81.45
+AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 10.25 96.35
+GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.69 2.29
+MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 22.46 85.71
+SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 50.88 48.24
+WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2462.11 4075.28
+WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 917.10 1210.33
+HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.44 5.92
+A.2.2 Monash
+Table 2: We present the mean MAE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
+llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
+Dataset
+australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 523.12 659.60
+bitcoin 1.75e18 5.33e18 5.33e18 9.9e18 1.10e18 3.62e18 6.66e18 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.97e18 5.32e18
+pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 58.35 170.88
+weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.17 2.36
+nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 4.09 8.26
+nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 14.33 16.71
+tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 93572.42 99456.05
+tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 10037.84 15845.10
+tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 2916.02 5636.83
+cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 896414.41 386526.37
+covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 555.24 2653.98
+fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 1823.16 2825.67
+traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.03
+traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.13 1.19
+saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 24.53 21.50
+us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 408.49 1152.67
+hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 19.34 24.07
+solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1184.01 1729.41
+Table 3: We present the mean msMAPE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
+llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
+Dataset
+australian electricity demand 15.30 22.07 22.18 13.70 44.22 28.75 8.76 8.35 9.22 12.77 7.74 8.07 8.68 17.27 22.07
+bitcoin 22.70 30.31 40.36 20.12 20.65 31.11 21.48 29.78 21.00 21.16 31.95 22.16 23.32 42.00 192.46
+pedestrian counts 53.02 121.39 122.08 119.76 148.48 138.58 40.29 45.54 39.30 36.10 54.15 34.22 36.16 50.73 121.35
+weather 34.05 50.85 56.19 58.06 51.47 57.98 106.01 59.12 38.17 36.46 50.87 40.13 35.35 50.85 36.81
+nn5 daily 47.43 35.38 21.93 21.11 21.49 25.91 30.20 24.04 23.30 23.75 28.47 22.65 23.18 23.46 48.09
+nn5 weekly 12.28 12.24 11.96 11.62 12.29 11.83 11.45 11.67 11.49 11.52 10.93 14.95 14.83 11.01 13.26
+tourism yearly 48.28 34.10 31.93 33.94 36.52 33.39 46.92 31.54 33.73 34.06 30.24 28.80 34.66 30.14 42.14
+tourism quarterly 25.25 27.41 15.37 17.16 15.07 16.58 15.86 16.53 16.20 15.29 14.45 15.56 16.97 19.92 31.68
+tourism monthly 33.53 36.39 19.89 21.20 19.02 19.73 21.11 21.10 20.11 18.35 20.42 18.92 19.74 22.77 40.40
+cif 2016 18.63 14.94 13.04 12.19 12.18 11.69 12.32 14.86 12.31 13.54 11.71 18.82 12.55 20.00 15.41
+covid deaths 20.65 15.35 15.57 8.71 8.64 9.26 18.34 15.40 18.58 34.18 32.36 14.50 40.71 15.89 134.79
+fred md 9.75 8.72 9.72 7.97 8.40 7.98 30.77 9.16 8.99 8.33 8.32 9.08 11.26 8.47 8.76
+traffic hourly 11.70 8.73 8.73 12.58 9.84 11.72 5.97 7.94 4.30 4.16 5.16 5.22 4.10 4.76 8.73
+traffic weekly 12.85 12.40 12.48 12.80 12.63 13.45 12.46 12.90 12.64 13.13 12.31 13.21 15.18 12.48 12.98
+saugeenday 58.53 35.99 35.97 37.34 67.50 37.55 45.32 35.55 39.32 40.22 56.03 37.02 56.62 43.31 35.99
+us births 4.48 11.77 5.82 3.81 4.05 5.17 5.75 4.23 5.55 4.13 4.17 4.88 4.36 3.94 11.35
+hospital 21.68 17.94 17.27 17.55 17.46 17.79 17.56 18.04 18.29 17.41 17.72 17.51 20.04 17.86 21.55
+solar weekly 37.33 24.59 24.76 19.05 22.93 17.87 21.65 29.35 21.52 15.00 24.05 32.50 12.26 24.20 32.80
+In Table 2 we present the actual MAE numbers that are behind the main Figure 2(a).
+It should be noted that though the scaled MAE metric accounts for scale variation across datasets, it does not account for within-dataset scale variations.
+Indeed, in datasets like cif 2016 and covid deaths, there are time-series with different order of magnitudes.
+The mean MAE for these datasets does not reflect average performance but only the performance of the higher scale time-series.
+Therefore in Table 3 we also present the normalized msMAPE metrics for TimesFM and all baselines.
+A.2.3 Informer
+We present the MAE on the last split of the test set for all dataset, horizon pairs considered in Table 4.
+Owing to expensive evaluations for llmtime, the results are reported on the last test window of the original test split, as done in [9].
+Table 4: MAE for ETT datasets for prediction horizons 96 and 192. Owing to expensive evaluations for llmtime, the results are reported on the last test window of the original test split.
+llmtime(ZS) PatchTST TimesFM(ZS) FEDFormer AutoFormer Informer
+Dataset
+ETTh1 (Horizon=96) 0.42 0.41 0.37 0.58 0.55 0.76
+ETTh1 (Horizon=192) 0.50 0.49 0.49 0.64 0.64 0.78
+ETTh2 (Horizon=96) 0.33 0.28 0.28 0.67 0.65 1.94
+ETTh2 (Horizon=192) 0.70 0.68 0.58 0.82 0.82 2.02
+ETTm1 (Horizon=96) 0.37 0.33 0.25 0.41 0.54 0.71
+ETTm1 (Horizon=192) 0.71 0.31 0.24 0.49 0.46 0.68
+ETTm2 (Horizon=96) 0.29 0.23 0.28 0.36 0.29 0.48
+ETTm2 (Horizon=192) 0.31 0.25 0.24 0.25 0.30 0.51
+A.3 More Details on Models
+We now present implementation details about TimesFM and other baselines.
+Monash Baselines.
+The raw metrics for the Monash baselines are directly taken from Tables 9 and 11 of the supplementary material of the original paper [8].
+For llmtime, we use the precomputed outputs provided by the authors of  [9].
+Darts Baselines. For all the Darts baselines we use the precomputed outputs provided by the authors of  [9]. For more details please see Section C.1 in that paper.
+TimesFM.
+For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128.
+The model dimension is set to 1280.
+We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $5e-4$.
+Informer Baselines.
+For FEDFormer [35], Autoformer [32], Informer [34] and PatchTST [19] we use the original hyperparameters and implementation.
+The results presented in the main paper are obtained on the last test window of length horizon length as stated in the llmtime [9] paper.
+We generate the llmtime predictions using the code provided by the authors  [^6] https://github.com/ngruver/llmtime/blob/main/experiments/run_monash.py but adapted to the ETT datasets.
+Note that as of January 2024, OpenAI has discontinued access to GPT-3, therefore we had to use the GPT-3.5-Turbo model.
+A.4 Synthetic Data
+We create the synthetic data to reflect common time series patterns using traditional statistical models. We start with four simple times series patterns:
+  • Piece-wise linear trends (I), where the number of the piece-wise linear components is randomly chosen between 2 and 8.
 […]
-!
-Model ETTh1 ETTh2 ETTm1 ETTm2 NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 $\pm$ 0.0004 0.380 $\pm$ 0.0002 0.245 $\pm$ 0.0004 0.161 $\pm$ 0.0004 0.571 $\pm$ 0.003 0.307 $\pm$ 0.001 0.180 $\pm$ 0.0008 0.114 $\pm$ 0.0007 TiDE(S) 0.663 $\pm$ 0.0001 0.374 $\pm$ 0.0001 0.245 $\pm$ 0.002 0.161 $\pm$ 0.001 0.588 $\pm$ 0.0005 0.320 $\pm$ 0.0005 0.186 $\pm$ 0.0005 0.120 $\pm$ 0.0001 N-BEATS(S) 0.687 $\pm$ 0.002 0.421 $\pm$ 0.002 0.235 $\pm$ 0.0002 0.152 $\pm$ 0.0001 0.608 $\pm$ 0.007 0.320 $\pm$ 0.004 0.179 $\pm$ 0.001 0.111 $\pm$ 0.0003 FEDFormer(S) 0.675 $\pm$ 0.003 0.411 $\pm$ 0.004 0.294 $\pm$ 0.010 0.205 $\pm$ 0.01 0.635 $\pm$ 0.007 0.381 $\pm$ 0.005 0.240 $\pm$ 0.006 0.152 $\pm$ 0.004 Model Wiki ILI TourismL - NRMSE WAPE NRMSE WAPE NRMSE WAPE - - PatchTST(S) 0.115 $\pm$ 0.0002 0.081 $\pm$ 0.0002 0.414 $\pm$ 0.005 0.132 $\pm$ 0.003 0.595 $\pm$ 0.012 0.204 $\pm$ 0.003 - - TiDE(S) 0.103 $\pm$ 0.0004 0.070 $\pm$ 0.0006 0.455 $\pm$ 0.004 0.131 $\pm$ 0.005 0.574 $\pm$ 0.001 0.194 $\pm$ 0.002 - - N-BEATS(S) 0.103 $\pm$ 0.0002 0.069 $\pm$ 0.0001 0.410 $\pm$ 0.003 0.137 $\pm$ 0.002 0.666 $\pm$ 0.002 0.213 $\pm$ 0.0003 - - FEDFormer(S) - - 0.441 $\pm$ 0.0007 0.164 $\pm$ 0.0003 1.080 $\pm$ 2 0.266 $\pm$ 0.0007 - -
-Table 7: NRMSE and WAPE confidence intervals (mean $\pm$ 1 standard error) of the supervised baselines.
+  • Seasonal patterns. In particular we create the sine (III) and the cosine (IV) waves of different random periods between 4 and 96 time-points and time delays.
+We then randomly enable / disable these four components (I) - (IV), generate their time series of length 1024 respectively, and sum them up using uniformly sampled random weights to create each times series in the synthetic datasets.
+We also choose to apply the trend multiplicatively 50% of the times the trend component is chosen.

### v4

@@
 Abstract
-Author names are listed in alphabetical order.1. Google Research, Correspond to senrajat@google.com
+Author names are listed in alphabetical order.
 Motivated by recent advances in large language models for Natural Language Processing (NLP), we design a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset.
-Our model is based on pretraining a patched-decoder style attention model on a large time-series corpus comprising both real-world and synthetic datasets.
+Our model is based on pretraining a decoder style attention model with input patching, using a large time-series corpus comprising both real-world and synthetic datasets.
 Experiments on a diverse set of previously unseen forecasting datasets suggests that the model can yield accurate zero-shot forecasts across different domains, forecasting horizons and temporal granularities.
-Machine Learning, ICML
-[…]
-[…]
 1 Introduction
@@
 In many of these domains, one of the most important use-cases of time-series data is forecasting.
-Time series forecasting is critical to several scientific and industrial applications, like retail supply chain optimization, energy and traffic prediction, and weather forecasting.
-In recent times, Deep learning models [25, 21] have emerged as a popular approach for forecasting rich, multivariate, time series data, often outperforming classical statistical approaches such as ARIMA or GARCH [3].
-In several forecasting competitions such as the M5 competition [17] and IARAI Traffic4cast contest [13], almost all the winning solutions are based on deep neural networks.
+Time-series forecasting is critical to several scientific and industrial applications, like retail supply chain optimization, energy and traffic prediction, and weather forecasting.
+In recent times, Deep learning models [30, 25] have emerged as a popular approach for forecasting rich, multivariate, time-series data, often outperforming classical statistical approaches such as ARIMA or GARCH [3].
+In several forecasting competitions such as the M5 competition [22] and IARAI Traffic4cast contest [15] deep network based solutions performed very well.
 At the same time, we are witnessing a rapid progress in the Natural Language Processing (NLP) domain on large foundation models for downstream NLP tasks.
@@
 This makes them very powerful tools that can be used for a variety of downstream tasks, often in a zero-shot learning mode.
-This motivates the question: “Can large pretrained models trained on massive amounts of time-series data learn temporal patterns that can be useful for time-series forecasting on previously unseen datasets?” In particular, can we design a time series foundation model that obtains good zero-shot out-of-the-box forecasting performance on previously-unseen datasets?
-Such a time series foundation model, if possible, would bring significant benefits for downstream forecasting users in terms of significantly reduced training data and compute requirements.
-It is not immediately obvious that such a foundation model for time series forecasting is possible.
+This motivates the question: “Can large pretrained models trained on massive amounts of time-series data learn temporal patterns that can be useful for time-series forecasting on previously unseen datasets?” In particular, can we design a time-series foundation model that obtains good zero-shot out-of-the-box forecasting performance ?
+Such a pretrained time-series foundation model, if possible, would bring significant benefits for downstream forecasting users in terms of no additional training burden and significantly reduced compute requirements.
+It is not immediately obvious that such a foundation model for time-series forecasting is possible.
 Unlike in NLP, there is no well defined vocabulary or grammar for time-series.
-Additionally, the model would need to support forecasting with varying history lengths (context) , prediction lengths (horizon) and time granularities.
-Furthermore, unlike the huge volume of public text data for pretraining language models, vast amounts of time series data is not readily available.
+Additionally, such a model would need to support forecasting with varying history lengths (context) , prediction lengths (horizon) and time granularities.
+Furthermore, unlike the huge volume of public text data for pretraining language models, vast amounts of time-series data is not readily available.
 In spite of these issues, we provide evidence to answer the above question in the affirmative.
-In particular, we design a single foundation model for time series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets).
+In particular, we design TimesFM, a single foundation model for time-series forecasting that, when applied to a variety of previously-unseen forecasting datasets across different domains, obtains close to state-of-the-art zero-shot accuracy (compared to the best supervised models trained individually for these datasets).
 Our model can work well across different forecasting history lengths, prediction lengths and time granularities at inference time.
-The key elements of our foundation model are twofold: 1) a large-scale time series corpus built using both synthetic and real-world (mostly using time series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/), that meets the volume and diversity of data needed for training our foundation model, and 2) a patched-decoder style attention architecture that can be efficiently pre-trained on this time series corpus.
-Compared to the latest large language models, our time series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (100B timepoints); yet we show that even at such scales it is possible to build a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time series data.
-Our work also suggests that unlike recent work [9] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a fraction of its costs.
+The key elements of our foundation model are twofold: 1) a large-scale time-series corpus built using both real-world (mostly time-series data from web search queries [^1] https://trends.google.com and Wikipedia page visits [^2] https://wikimedia.org/api/rest_v1/) and synthetic data, which meets the volume and diversity of data needed for training our foundation model, and 2) a decoder style attention architecture with input patching, that can be efficiently pre-trained on this time-series corpus.
+Compared to the latest large language models, our time-series foundation model is much smaller in both parameter size (200M parameters) and pretraining data size (O(100B) timepoints); yet we show that even at such scales, it is possible to pretrain a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised approaches on a diverse set of time-series data.
+Our work also suggests that unlike recent work [11] that recommends Large Language Models such as GPT-3 and LLama-2 as out-of-the-box zero-shot forecasters, foundation models trained from scratch exclusively on time-series data can obtain much better zero-shot performance at a tiny fraction of its costs.
 2 Related Work
-In the last decade, deep learning models [25, 21] have emerged as powerful contenders in forecasting time-series in the presence of large training datasets and have been shown to outperform traditional statistical methods such as ARIMA, Exponential smoothing [18].
+In the last decade, deep learning models [30, 25] have emerged as powerful contenders in forecasting time-series in the presence of large training datasets and have been shown to outperform traditional statistical methods such as ARIMA and Exponential smoothing [20].
 Forecasting models can be categorized broadly into: (i) Local univariate models that include traditional methods like ARIMA, exponential smoothing [20] and non-autoregressive models like Prophet [32].
@@
 All the works cited above have primarily been applied in the supervised setting with the notable exception of PatchTST [24] and N-BEATS [25].
-PatchTST has a section on dataset to dataset transfer learning in the semi-supervised setting.
-The patching in our decoder-only model is inspired by [19].
+PatchTST has a section on dataset-to-dataset transfer learning in the semi-supervised setting.
 [26] also show that the N-BEATS architecture lends itself to transfer learn between various source-target dataset pairs.
 However, none of these works aim to train a single foundation model that can work on a plethora of datasets.
-For a more in-depth discussion about transfer learning in time-series we refer the reader to the survey in [16].
-There has been some recent work on re-using or fine-tuning large language models for time-series forecasting.
+For an in-depth discussion about transfer learning in time-series we refer the reader to the survey in [21].
+There has been some very recent work on re-using or fine-tuning large language models for time-series forecasting.
 In particular,  [11] benchmarks pretrained LLMs like GPT-3 and LLaMA-2 for zero-shot forecasting performance.
 As we show later, our model obtains much superior zero-shot performance at a tiny fraction of these model sizes.
-Zhou et al.
 [42] and [6] show how to fine-tune a GPT-2 [29] backbone model for time-series forecasting tasks.
-Both the works have a section on zero-shot forecasting on a target dataset after having trained on a source dataset.
-With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on a target dataset, and not on pretraining a single foundation model that shows out-of-the box zero-shot performance on a variety of datasets.
+With the exception of a transfer-learning study (forecasting on a target dataset after having trained on a source dataset), these papers mostly focus on fine-tuning a pretrained model on target datasets, and not on pretraining a single foundation model with good out-of-the box zero-shot performance on a variety of datasets.
 To the best of our knowledge, the very recent work in TimeGPT-1 [12] is the only other parallel work on a zero-shot foundation model for time-series forecasting.
-However the model is currently not public access, and several model details and the benchmark dataset have not been revealed.
+However the model is not public access, and several model details and the benchmark dataset have not been revealed.
 3 Problem Definition
@@
 Let the context be denoted by $\mathbf{y}_{1:L}:=\{y_{1},\cdots,y_{L}\}$ where we follow a numpy-like notation for indices.
-Similarly the actual values in the horizon is denoted by $\mathbf{y}_{L+1:L+H}$.
+Similarly the actual values in the horizon are denoted by $\mathbf{y}_{L+1:L+H}$.
 Note that since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time.
-However, the datetime column is ubiquitous in all time-series data, so we can optionally have date derived features like day or the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$.
-Such features could be available for forecasting in both the context and horizon, represented as $\mathbf{x}_{1:L+H}$.
-The task is then to learn a capable foundation model that can map any time-series context to horizon, given by
-$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L},{\bm{x}}_{1:L+H}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$ (1)
-The accuracy of the prediction will be measured by a metric that quantifies their closeness to the actual values.
-For instance, if the metric is Mean Squared Error (MSE), then the goodness of fit is measured by,
-$\mathrm{MSE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$
-$=\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{2}^{2}.$ (2)
+The task is then to learn a foundation model that can map any time-series context to horizon,
+$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$ (1)
+The accuracy of the prediction can be measured by a metric that quantifies their closeness to the actual values, for instance, Mean Absolute Error (MAE) defined in Equation 6.
 4 Model Architecture
@@
 Inspired by the success of patch based modeling in the recent long horizon forecasting work [24] we also choose to break down the time-series into patches during training.
-A patch of a time-series is a natural analogues for a token in language models and have been shown to improve performance.
+A patch of a time-series is a natural analogue for a token in language models and patching has been shown to improve performance.
 Moreover this improves inference speed as the number of tokens being fed into the transformer is reduced by a factor of the patch length.
@@
 A key difference between our architecture and PatchTST [24] is that our model is trained in decoder-only mode [18].
-In other words, given a sequence of input patches the model is optimized to predict the next patch as a function of all past patches.
-Similar to LLM’s this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
+In other words, given a sequence of input patches, the model is optimized to predict the next patch as a function of all past patches.
+Similar to LLMs this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.
 Longer output patches.
-In LLM’s the output is always generated in an auto-regressive fashion one token at a time.
-However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy that multi-step auto-regressive decoding [33].
-But the latter is not possible when the horizon length is not known a prior like in the case of zero-shot forecasting, which is our primary goal.
-We propose a middle ground by allowing the output patches for prediction to be be longer than the input patches.
+In LLMs the output is always generated in an auto-regressive fashion one token at a time.
+However, in long-horizon forecasting it has been observed that directly predicting the full horizon yields better accuracy than multi-step auto-regressive decoding [40].
+But this is not possible when the horizon length is not known apriori, as in the case of zero-shot forecasting which is our primary goal.
+We propose a middle ground by allowing our output patches for prediction to be longer than the input patches.
 As an example, suppose the input patch length is 32 and output patch length is 128.
@@
 The model will first generate the future predictions for time-steps 257 to 384, then condition on the initial 256 length input plus the generated output to generate time-steps 385 to 512.
-On the other hand, if in our model the output patch length was equal to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above.
+On the other hand, if in a model the output patch length was fixed to the input patch length of 32, then for the same task we would have to go through 8 auto-regressive generation steps instead of just the 2 above.
+However, there is a trade-off.
+If the output patch length is too long, then it is difficult to handle time-series whose lengths are less than the output patch length for instance monthly, yearly time-series in our pretraining data.
 Patch Masking.
@@
 Parts of patches as well as entire patches from the beginning of the context window can be masked in a data batch.
-We employ a specific random masking strategy during training that helps to model see any all possible context lengths starting from 1 to a maximum context length.
-Now that we have mentioned the guiding principles, we next formally describe each component of our architecture that is illustrated in Figure 1.
-Figure 1: We provide an illustration of our model architecture during training where we show a input time-series of a certain length that can be broken down into input patches.
-Each patch along with (optional) time-features is processed into a vector by a residual block (as defined in the model definition) to the model dimension of the transformer layers.
+We employ a specific random masking strategy (described later) during training that helps the model see all possible context lengths starting from 1 to a maximum context length.
+Now that we have mentioned the guiding principles, we next formally describe each component of our model architecture (illustrated in Figure 1), which we name as TimesFM (Time-series Foundation Model).
+Figure 1: We provide an illustration of the TimesFM model architecture during training, where we show a input time-series of a specific length that can be broken down into input patches.
+Each patch along is processed into a vector by a residual block (as defined in the model definition) to the model dimension of the transformer layers.
 The vector is then added to positional encodings and fed into $n_{l}$ stacked transformer layers.
 SA refers to self-attention (note that we use multi-head causal attention) and FFN is the fully connected layer in the transformer.
-The output tokens are then mapped through a residual block to an output of size output_patch_len which is the forecast for the time-period following the last input patch seen by the model so far.
+The output tokens are then mapped through a residual block to an output of size output_patch_len, which is the forecast for the time window following the last input patch seen by the model so far.
 Input Layers.
@@
 Along with the input, we also supply a binary padding mask $\mathbf{m}_{1:L}$ where $1$ denotes that the corresponding input in $\mathbf{y}_{1:L}$ should be ignored and vice-versa.
-The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection as defined in [6].
+The Residual Block is essentially a Multi-layer Perceptron (MLP) block with one hidden layer with a skip connection, similar to that defined in [7].
 In other words, the inputs $\mathbf{y}_{1:L}$ are broken down into patches of size input_patch_len ($p$).
@@
 The main hyperparameters are model_dim which is equal to the dimension of the input tokens $\mathbf{t}_{j}$’s and number of heads (num_heads).
-We set the hidden size of the FFN’s to be equal to model_dim as well.
+We set the hidden size of the FFNs to be equal to model_dim as well.
 We use causal attention that is each output token can only attend to input tokens that come before it in the sequence (including the corresponding input token).
 This can be described by the equation
-$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$
-$\forall j\in[N].$ (4)
-In Equation (4), $\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked.
+$,\displaystyle\mathbf{o}_{j}=\texttt{StackedTransformer}((\mathbf{t}_{1},\dot{m}_{1}),$$\cdots,(\mathbf{t}_{j},\dot{m}_{j})),$ (3)
+for all $j\in[N]$.
+$\dot{m}_{j}$ is the masking indicator for the $j$-th token defined as $\min\{\mathbf{m}_{p(j-1)+1:pj}\}$ i.e if a patch has any non-masked time-point the corresponding token marked as not being masked.
 All patches that are masked out completely are not attended to by the causal self attention.
@@
 In this work, we focus on point forecasting.
-Therefore we can use a point forecasting loss during training like MSE as defined in Equation (2).
+Therefore we can use a point forecasting loss during training like Mean Squared Error (MSE).
 The loss that is minimized during training can be expressed as,
@@
 Training.
-We train the model with standard mini-batch gradient descent in decoder only fashion, that goes through all windows for a time-series and across time-series.
+We train the model with standard mini-batch gradient descent in decoder-only fashion, that goes through all windows for a time-series and across time-series.
 The only non-standard part is the way we sample the mask during training.
 For each time-series in the batch, we sample a random number $r$ between 00 and $p-1$.
-Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask our a fraction of the first input patch.
+Then we set the $\mathbf{m}_{1:r}=1$ and the rest as zero i.e we mask out a fraction of the first input patch.
 However, this is sufficient to cover all input context lengths from 1 to the maximum training context length.
-Let us say that the maximum context length is 512 and $p=32$.
-Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized predict after seeing $28=32-4$ time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing $28+32$ time-points, and so on.
-When this argument is repeated for all such $r$’s, we can see that the model has seen all possible context lengths from $1$ to $512$.
+We explain this using an example below:
+Suppose the maximum context length is 512 and $p=32$.
+Then if $r=4$, the output prediction after seeing the first patch (from $\mathbf{o}_{1}$) is optimized to predict after seeing $28=32-4$ time-points, the output of the next patch (from $\mathbf{o}_{2}$) is optimized to predict after seeing $28+32$ time-points, and so on.
+When this argument is repeated for all such $r$’s, the model has seen all possible context lengths till $512$.
 Inference.
@@
 Then, we can use the concatenated vector $\tilde{\mathbf{y}}_{1:L+h}=[\mathbf{y}_{1:L};\hat{\mathbf{y}}_{L+1:L+h}]$ as an input to the network to generate the next output patch prediction $\hat{\mathbf{y}}_{L+h+1:L+2h}$ and so on.
-If $L$ is not a multiple of $p$ we can append some zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.
-We name our Time-series Foundation Model as TimesFM.
+If $L$ is not a multiple of $p$, we simply append zeros to make it a multiple of $p$ and mark the corresponding entries in the mask as 1.
 5 Pretraining Details
-We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model.
+We would like our pretraining corpus to include large volumes of temporal data representing a variety of domains, trend and seasonality patterns and time granularities that ideally capture the forecasting use-cases which we are interested in serving by the deployed model.
 It is challenging to find a large time-series dataset that meets the volume and diversity of data needed for training our foundation model.
-The bulk of real data used to train our models come from Google trends  [^3] https://trends.google.com and Wiki Pageview statistics  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics.
+We address this problem by sourcing the bulk of data used to train our models from three major sources: Google trends, Wiki Pageview statistics and synthetic time-series.
 In summary the main data sources are:
 Google Trends.
-Google Trends captures search interest over time for millions of queries.
+Google Trends  [^3] https://trends.google.com captures search interest over time for millions of queries.
 We choose around 22k head queries based on their search interest over 15 years from 2007 to 2022.
-We use the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset.
+Beyond these head queries the time-series become more than 50% sparse.
+We download the search interest over time for these queries in hourly, daily, weekly and monthly granularities to form our dataset.
 The date ranges are Jan.
@@
 2021 for the other granularities.
+The trends datasets amounts to roughly 0.5B time-points.
 Wiki Pageviews.
-Wiki Pageviews captures the hourly views of all Wikimedia pages.
+Wiki Pageviews  [^4] https://en.wikipedia.org/wiki/Wikipedia:Pageview_statistics captures the hourly views of all Wikimedia pages.
 We download all pageview data from Jan.
 2012 to Nov.
-2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out the resulting time-series with excessive zeros.
-The final corpus contains O(100B) time points.
-Other real data sources.
+2023, clean and aggregate the views by page into hourly, daily, weekly and monthly granularities, and filter out pageview time-series with excessive zeros.
+The final corpus contains roughly 300B time-points.
+Synthetic Data.
+Another major component of our pretraining data is of synthetic origin.
+We create generators for ARMA [20] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions.
+A synthetic time-series can be an additive combination of one or more of these processes.
+We create 3M synthetic time-series each of length 2048 time-points.
+More details about our synthetic data generation are presented in Appendix A.8.
+Other real-world data sources.
 Along with the wiki and trends data, we also add time-series from several other publicly available datasets to our pretraining corpus.
-We add in all the granularities of the M4 dataset [17] and the hourly (and 15 minute) Electricity and hourly Traffic datasets (see  [34]).
-We also add the 10 minute granularity Weather dataset used for evaluations in [34].
+We add all the granularities of the M4 dataset [22], the hourly and 15 minute Electricity and the hourly Traffic datasets (see  [44]).
+We also add the 10-minute granularity Weather dataset used for evaluations in [44].
 M4 has a good mix of granularities with around 100k time-series in total.
 Traffic and Electricity are large long-term forecasting datasets with $>$ 800 and $>$ 300 time-series each having tens of thousands of time-points.
-In addition, we add all the 15 min granularity traffic time series from [29].
-Synthetic.
-Another major component of our pretraining data is of synthetic origin.
-We create generators for ARMA [18] processes, seasonal patterns (mixture of sines and cosines of different frequencies), trends (linear, exponential with a few change-points) and step functions.
-A synthetic time-series can be an additive combination of 1 or more of these processes.
-We create 3M synthetic time-series each of length 2048 time-points.
-More details are presented in Appendix A.4.
-Dataset Mixing.
-We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities.
-The training mixture is 50% real data and 50% synthetic, with the real data mixture providing equal weights to all hourly + sub-hourly, daily, weekly, and monthly datasets.
+In addition, we add all the 15 min granularity traffic time-series from [36].
+Table 1: Composition of TimesFM pretraining dataset.
+Dataset Granularity # Time series # Time points
+Synthetic 3,000,000 6,144,000,000
+Electricity Hourly 321 8,443,584
+Traffic Hourly 862 15,122,928
+Weather [44] 10 Min 42 2,213,232
+Favorita Sales Daily 111,840 139,179,538
+LibCity [36] 15 Min 6,159 34,253,622
+M4 hourly Hourly 414 353,500
+M4 daily Daily 4,227 9,964,658
+M4 monthly Monthly 48,000 10,382,411
+M4 quarterly Quarterly 24,000 2,214,108
+M4 yearly Yearly 22,739 840,644
+Wiki hourly Hourly 5,608,693 239,110,787,496
+Wiki daily Daily 68,448,204 115,143,501,240
+Wiki weekly Weekly 66,579,850 16,414,251,948
+Wiki monthly Monthly 63,151,306 3,789,760,907
+Trends hourly Hourly 22,435 393,043,680
+Trends daily Daily 22,435 122,921,365
+Trends weekly Weekly 22,435 16,585,438
+Trends monthly Monthly 22,435 3,821,760
+Dataset Mixing and Training.
+We train on a mixture distribution over these datasets that aims to give sufficient weight to all granularities and datasets.
+The training loader samples 80% real data and 20% synthetic, with the real data mixture providing equal weights to the groups: hourly + sub-hourly, daily, weekly, and monthly datasets.
 We train with a maximum context length of 512 whenever the length of the time-series allows that.
-For weekly granularity we do not have sufficiently long time-series therefore a maximum context length of 256 is used.
+For weekly granularity we do not have sufficiently long time-series; therefore a maximum context length of 256 is used.
 For the same reason, a maximum context length of 64 is used while training on $\geq$ monthly granularity data.
+We also use only the standard normalization part of reversible instance normalization [16] – i.e, the context of each time-series is scaled by the context mean and standard deviation of the first input patch in the context.
 (a) Monash Archive [9]
@@
 In all figures, the lower the metric the better and the error bars represent one standard error.
-Note that among the baselines onlye TimesFM and llmtime are zero-shot.
+Note that among the baselines only TimesFM and llmtime are zero-shot.
 In (a) we report results on the Monash datasets.
-Since the datasets have different scales, we average the MAE’s scaled by the MAE of a naive baseline.
-We can see that TimesFM is among the top-3 models.
+Since the datasets have different scales, we take the Geometric Mean (GM) of the MAE’s scaled by the MAE of a naive baseline.
+We can see that TimesFM is the top model.
 In (b), we report the similarly scaled MAE on the Darts benchmarks.
-TimesFM is within significance of the top of the best performing method which is ARIMA in this case.
+TimesFM is within significance of the best performing methods which are ARIMA and llmtime in this case.
 Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones.
 Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total.
-TimesFM and PatchTST are the best performing models in this case.
+TimesFM and PatchTST are the best performing models
 6 Empirical Results
@@
 To benchmark our model’s performance, we choose three groups of commonly used forecasting datasets that cover various domains, sizes, granularities, and horizon lengths: Darts [14], Monash [9] and Informer datasets [44], to test the generalization power of our foundation model against other baselines.
-In all cases, we report performance on the official metrics and scalings of the datasets. We present a summary of the results below - more detailed experimental results can be found in Appendix A.2.
+In all cases, we report performance on the official metrics and scalings of the datasets, using either their standard test splits or common test splits in other literature.
+We present a summary of the results below - more details can be found in Appendix A.5.
+We provide the hyper-parameters and other details about our model in Appendix A.6.
 Monash [9].
@@
 Following llmtime [11] we start from the Monash Huggingface repository  [^5] https://huggingface.co/datasets/monash_tsf and filter out the datasets that contain missing values.
-This leaves us with 18 datasets.
+This leaves us with 18 datasets which we specify in Appendix A.5.2.
 Out of the four official metrics, following prior work [11], we report our performance in terms of mean MAE (see Appendix A.2).
@@
 The scaled aggregation was also used in [11].
+In Figure 2(a), we use the Geometric Mean (GM) for averaging since it is more robust for normalized metrics [8].
+We also report the Arithmetic Mean based aggregated metrics in Figure 4 in the appendix.
 The mean scaled MAE across all datasets is plotted in Figure 2(a) along with standard error bars.
@@
 Note that the zero-shot models are marked as (Zero-Shot).
-TimesFM is among the top 3 models even though we never trained on these datasets.
-It outperforms some deep supervised models like DeepAR and FFNN, and improves on llmtime’s performance by more than 10%.
+TimesFM is the top model even though we never trained on these datasets.
+It is slightly better but within significance of N-BEATS but outperforms deep supervised models like DeepAR [30], and improves on llmtime’s performance by more than 25%.
 Darts [14].
@@
 As before, we also report zero-shot forecasting results from llmtime [11] using GPT-3 [29].
-Other supervised baselines implemented in [9] like SM-GP [31] and seasonal ARIMA [18] are also added.
+Other supervised baselines in [11] like SM-GP [35] and ARIMA [20] are also added.
 We report the official metric for this dataset group that is MAE for each individual dataset in Appendix A.5.
 In Figure 2(b), we present the average scaled MAE across all 8 datasets, as we did for the Monash datasets.
-TimesFM is within statistical significance of the best model that is seasonal ARIMA in this case.
+TimesFM is within statistical significance of the best models that is llmtime and seasonal ARIMA in this case.
 Note that since there are only 8 individual time-series in this dataset group, the standard errors are not sharp and therefore does not provide a clear ordering among the models.
+Also, note that for ARIMA, the seasonality needs to be encoded correctly in the parameters for the best results, which needed manual tuning.
+Further, since these datasets are used in numerous time series blog posts for illustrative purposes, data contamination for llmtime cannot be ruled out.
 Informer [44].
@@
 Note that the long horizon baselines usually report rolling validation results on the test set which would amount to millions of tokens for evaluating llmtime [11] and would be too expensive.
-Therefore, following the latter paper, we compare all methods on the last test window.
+Therefore, following llmtime, we compare all methods on the last test window.
 Also, it is reasonable to directly average the MAE for these datasets since the results are reported on standard normalized dataset (using the statistics of the training portion).
@@
 The other long horizon methods are quite a bit worse even though they have been trained these datasets.
-llmtime(ZS) is better than FEDFormer but worse than PatchTST in a statistically significant way.
+llmtime is better than FEDFormer but worse than PatchTST with statistical significance.
+We present visual examples of our forecasts along with baselines in Appendix A.9.
 6.2 Ablation
+(a) Scaled MAE (GM) on Monash datasets as a function of FLOPS across three model sizes 17M, 70M and 200M. The first 4 points are from 17M and 70M checkpoints while the last 3 are from 200M.
+(b) Ablation with respect to output patch length for the task of predicting 512 steps into the future on ETT datasets on the original test set in [44]. We report the average across all 4 ETT datasets.
+(c) Scaled MAE (GM) for our 70M models on Monash datasets for different input patch lengths. We also plot error bars denoting one standard error.
+(d) Average scaled MAE for Monash on the left and average MAE on ETT datasets on the right. We compare the performance of 200M model with and without the synthetic data.
+Figure 3: Ablation studies with respect to various design choices.
 Next, we perform several ablation studies that inform the design decisions we made for our model architecture.
@@
 Performance curves with respect to number of parameters in a model have been a keenly studied area in the context of LLMs.
-Kaplan et al.
 [17] established a power law like relationship between the number of parameters in a language model and its downstream performance i.e the more the number of paramaters the better the performance.
-However,  Hoffmann et al.
-[11] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset.
-We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pretraining dataset and till the similar number of iterations.
-We provide the mean scaled MAE results on the Monash datasets in Figure 3.
-It can be clearly seen that the errors decrease monotonically with the number of parameters, thus pointing to a similar observation as [12], but in the context of time-series.
-Figure 3: Average scaled MAE on Monash datasets for three different TimesFM model sizes. We can see that the performance improves with increasing model size.
+However,  [13] established a more nuanced scaling law that lays down methods to train compute optimal models based on the number of tokens available in a training dataset.
+We perform a preliminary scaling study where we train three TimesFM models of sizes 17M, 70M and 200M parameters, using the same pre-training dataset till 1.5M iterations with a global batch-size of 4096.
+Then we collect checkpoints that represent varying number of FLOPS (Floating Point OPerationS) across the different model runs.
+Then we plot the performance on Scaled MAE (GM) on Monash as a function of FLOPS, in Figure 3(a).
+This is now a standard way to perform scaling studies in LLMs (see recent work like [10]).
+It can be clearly seen that the errors decrease monotonically with the number of FLOPS (in log scale).
+All experiments were performed on a TPUv5e [^6] https://cloud.google.com/tpu/docs/v5e-training setup with 16 tensor-cores.
+For the 200M model it takes 2 days to complete 1.5M iterations on our setup.
 Autoregressive Decoding.
 In recent long-term forecasting works [40, 24, 7] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks.
-For a foundation model the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons.
-However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps.
+For a foundation model, the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons.
+However, as mentioned earlier, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps.
 This was one of the key decisions in the design of TimesFM, that is quite different from LLMs.
-In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets.
-In Figure 4, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128.
-The former has to perform 16 autoregressive steps while the latter has to do only 4.
-It can be clearly seen that having a larger output_patch_len helps in this case.
-Figure 4: Ablation with respect to output patch length for the task of predicting 512 steps into the future, in the same setting as that of Section 6.1.
-We report the average across all ETT 4 datasets.
+In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets on the original rolling validation task of the ETT test sets [44].
+In Figure 3(b), we present results from models with output_patch_len varying from 8 to 128.
+We see a monotonic decrease in average MAE with output_patch_len.
 Input Patch Length.
@@
 We have typically seen that increasing its value from 8 to 32 increases performance but having too high a input_patch_len is impractical since that makes the model shift from decoder only training more towards encoder-decoder style training.
-Note that in the ”Training” paragraph of Section 4, we describe the mask sampling strategy to support any context length.
+Note that in the "Training" paragraph of Section 4, we describe the mask sampling strategy to support any context length.
 If in the extreme case $p$ is set the maximum context length we have to individually sample all possible context windows from 1 to maximum context length, which would be required for encoder-decoder style of training.
-In Figure 5 we show the mean scaled MAE TimesFM(ZS) - 70M model with input_patch_len=8 on Monash datasets, which is clearly worse than our original model that uses input_patch_len=32.
-Figure 5: Average scaled MAE for our 70M models on Monash datasets for two different input patch lengths. We can see that the performance improves with increasing patch size.
-6.3 Visualizing Forecast
-We conduct a visual inspection of the forecasts generated by TimesFM(ZS) on the benchmark datasets.
-Our inspection in particular emphasizes time-series exhibiting both trend and seasonal components readily interpretable by humans.
-As illustrated in Figure 6, TimesFM(ZS) effectively captures these subtle characteristics within both the trend and seasonal patterns of the depicted time-series.
-Figure 6: Forecasts visualized on Darts and Monash datasets. Plots on the right zoom in the forecasting parts from the left.
+In Figure 3(c), we show the mean scaled MAE (GM) TimesFM(ZS) - 70M model on Monash with input_patch_len varying from 8 to 128.
+Note that both models have been trained to about 1.5M steps even though the p=8 model is three times slower to train.
+We can see that $p=16,32$ marks the best performance, with the error increasing towards either end.
+Note that $p=32$ model is almost twice as fast to train compared to $p=16$ and thus constitutes a prudent choice.
+Dataset Ablation.
+Next we showcase the need for synthetic data.
+Intuitively, the majority of our real datasets have commonly found granularities like hourly, daily etc which have specific periodic patterns like 24 time-point period for hourly data.
+This can make the model not generalize well to underrepresented frequencies.
+We train a 200M model with no synthetic data added in the mix and showcase the performance on Monash and ETT datasets in Figure 3.
+It can be seen that there is a performance drop on Monash because many of the datasets in Monash have under-represented granularities like quarterly, yearly or 10 minutes etc.
+Perhaps even more compelling is the comparison on ETT datasets.
+We can see that there is almost no difference between the two models on the hourly ETTh datasets which has a well represented granularity.
+However, for the 15min ETTm datasets the model with synthetic data performs quite a bit better.
+We provide a finetuning study in the same setting as [42] in Appendix A.3, where our model performs better than all baselines on all the reported datasets.
+This shows the utility of our model on downstream tasks.
 7 Conclusion
-In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time series data.
-This model is trained on real-world and synthetic datasets comprising around 100B timepoints using a patched-decoder style attention architecture with around 200M parameters.
-In future work, we plan to delve into a more theoretical understanding of how such a time-series foundation model can obtain good performance for out-of-distribution data, and also investigate the fine-tuning/few-shot performance of this model.
+In this paper, we presented TimesFM, a practical foundation model for forecasting whose zero-shot performance comes close to the accuracy of fully-supervised forecasting models on a diverse set of time-series data.
+This model is pretrained on real-world and synthetic datasets comprising O(100B) timepoints.
+We discuss limitations and future work in more detail in Appendix A.1.
+8 Impact Statement
+This paper shows that it is possible to train a single pretrained model that has phenomenal zero-shot performance on a variety of forecasting tasks, thus opening up exciting possibilities for downstream applications.
+Therefore it is crucial to discuss ethical and societal considerations of using such a model and how some of the related concerns can be mitigated.
+Data Privacy.
+Note that most of our data sources are publicly available and are aggregated i.e no individual user activity constitutes a time-point.
+Further the Google Trends data is differentially private.
+Bias.
+Biases can creep into a foundation model through a variety of sources especially through data.
+The model might perpetuate these biases in forecasts, leading to unfair outcomes.
+Biased forecasts can have real-world consequences.
+For example, a biased forecast of crime rates in a neighborhood could lead to increased police presence, disproportionately impacting certain communities.
+Note that our model is not trained with any covariates so some of these sensitivities are reduced but cannot be ruled out.
+In this regard, we believe that it is best to release the exact details of the datasets used for training and therefore we have summarized our data sources in Table 1.
+Moreover we plan to do a open weights release of our model so that the community can analyze the model for downstream tasks.
+We will ensure that we release a good model card [23].
+This will also aid in finetuning the model with more diverse data-sources.
+Training cost.
+Our largest model has 200M parameters which is much smaller compared to even smaller scale LLMs.
+Our data volume is quite large but still smaller compared to SOTA LLMs.
+We have revealed the exact computation requirements for the models i.e 16 core TPUv5e for 2 days.
+Note that experimentation and trial runs of course cost more than the final model run.
+Therefore in the interest of equitability we would like to release the weights of our model in a responsible manner.
+Lastly we would like to note that similar to LLMs there could be inputs on which the model does not perform well or hallucinates.
+Therefore in many critical use cases it might be recommended to use the model in a human-in-the-loop fashion or alternatively do a wide range of testing or finetuning.
 Appendix A Appendix
-A.1 Metrics
+A.1 Limitations and Future Work
+Our work shows that we can train a 200M parameter pretrained forecasting model that has impressive zero-shot performance on a variety of real world forecasting benchmarks with different context and horizon lengths.
+In this section we would like to discuss limitations and future work.
+Prompt Tuning.
+In LLMs it is well known that prompt tuning techniques like chain-of-thought [38] can drastically improve performance in cases where the model is inaccurate with simple prompts.
+Such techniques are less clear for time-series foundation model.
+We can tune simple hyper-parameters like context length as the moment.
+However, with probabilistic forecasting we might be able to output different statistics as well as come up with techniques that align more with user’s expectations while not decreasing likelihood.
+Probabilistic Forecasting.
+It should be straightforward to train with probabilistic loss functions in our framework as detailed in the "Loss Function" part of Section 4.
+However, being one of the first works of building a single foundation model for forecasting, this was not our main focus and is left to future explorations.
+Note that as mentioned before we plan to release our model weights and after that such loss functions [30, 1] can be added during finetuning.
+Covariate handling.
+Currently the model is not pretrained with covariates as one of the key challenges is finding large volumes of pretrained data with meaningful covariates (apart from date features).
+We also need methods to have a joint universal representation of covariates.
+Currently there are two simple techniques we can think of for handling covariates (i) In a zero-shot setting at inference time we can predict in-context and linearly regress the residual on covaraites.
+Then our model + the residual model can be used for forecasting in the horizon.
+(ii) during finetuning it is straightforward to handle covariates by adding them as inputs to the input and output residual blocks.
+Categorical variables can be added as embeddings.
+More finetuning studies.
+We perform a fintuning study in Appendix A.3 following a prior work.
+However, a more in depth study that involves finetuning in the presence of covariates would be beneficial.
+This being one of the first works of building a single foundation model for forecasting, this was not our main focus and is left to future explorations.
+Ideas in recent work such as [4] could be useful in this regard.
+Other architectures.
+Given the cost of training foundation models we did not perform much hyper-parameter tuning in our pretraining, while following some well established best practices for training transformers.
+In a similar vein, it would also be interesting to try out alternatives like the exciting directions of all MLP structures like [4] or efficient linear state space models like Mamba [10] (and references there in).
+Interpretability.
+Deep foundation models trained on a huge corpuses of data could be inherently less interpretable compared to statistical methods like ARIMA, ETS [3].
+In this regard methods like LOCO, SHAP (see [34] and references there in) could be used to some extent to attribute feature importances to different lags in the context supplied to the model.
+However, this does not solve the problem to a full extent and one of the best things to do would be to open source a version of the model with a proper model card.
+[23].
+A.2 Metrics
 The metrics that are used for reporting results in this paper are:
@@
 We did not need to do that for the Informer datasets since on these datasets metrics are usually reported on standard normalized data [24].
-A.2 Additional Empirical Results
-In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1
-A.2.1 Darts
+A.3 Finetuning study on ETT
+In this section, we test whether TimesFM can be fintuned on a small fraction of a dataset to provide even better performance.
+We follow the same protocol as in GPT4TS [42] (see Table 13 in their paper).
+[42] finetune GPT2 input and output blocks on long-term forecasting benchmarks on 10% of the of the original datasets and compare it against models trained from scratch on the same data.
+Then the models are evaluated on the original test set task of [44].
+We also tune the input and output residual blocks on 10% of the training set and present the results in Table 2.
+We can see that our model performs the best by a large margin.
+In ETTh1, ETTh2, ETTm1 our finetuned model is better than 18%, 3% and 12% better than GPT4TS, respectively.
+In fact we can see our 10% finetuned model’s performances are comparable or better than that of most baselines trained on the whole training dataset as reported in Table 14 of  [42].
+This shows that the inductive biases encoded in our model weights by finetuning on a large time-series corpus are better for downstream forecasting task than an off the shelf language model like GPT2, even though our model is orders of magnitude smaller.
+Table 2: MAE of different methods on ETT datasets. All methods use 10% of the original training set for training or finetuning. The baseline numbers are from Table 13 in [44].
+Dataset TimesFM(FT) GPT4TS(FT) DLinear PatchTST TimeNet FEDFormer Autoformer
+ETTh1 96 0.398 0.456 0.495 0.485 0.628 0.499 0.552
+192 0.424 0.516 0.538 0.524 0.593 0.555 0.598
+336 0.436 0.535 0.622 0.550 0.648 0.574 0.619
+720 0.445 0.591 0.743 0.610 0.641 0.614 0.616
+Avg 0.426 0.525 0.600 0.542 0.628 0.561 0.596
+ETTh2 96 0.356 0.374 0.411 0.389 0.409 0.416 0.451
+192 0.400 0.411 0.519 0.414 0.467 0.474 0.477
+336 0.428 0.433 0.572 0.441 0.494 0.501 0.543
+720 0.457 0.464 0.648 0.480 0.491 0.509 0.523
+Avg 0.410 0.421 0.538 0.431 0.465 0.475 0.499
+ETTm1 96 0.345 0.404 0.392 0.419 0.501 0.518 0.614
+192 0.374 0.423 0.412 0.434 0.528 0.546 0.592
+336 0.397 0.439 0.434 0.454 0.568 0.775 0.677
+720 0.436 0.498 0.477 0.556 0.549 0.579 0.630
+Avg 0.388 0.441 0.429 0.466 0.537 0.605 0.628
+ETTm2 96 0.263 0.269 0.303 0.274 0.285 0.399 0.454
+192 0.309 0.309 0.345 0.317 0.323 0.379 0.691
+336 0.349 0.346 0.385 0.353 0.353 0.559 1.407
+720 0.415 0.417 0.440 0.427 0.449 0.614 1.166
+Avg 0.334 0.335 0.368 0.343 0.353 0.488 0.930
+A.4 Pretraining PatchTST
+Since TimesFM applies a similar patching strategy as PatchTST [24], for an ablation study we use the same data loader and pretrain a PatchTST model of 200M parameters to the same number of FLOPS as the final 200M TimesFM model.
+We denote it as PatchTST(ZS).
+The two models share the same hyperparameters of the transformer stack.
+For PatchTST(ZS) we use the same input patch length = 32, and a stride of length half of input patch size (i.e.
+stride = 16) as done in the original PatchTST paper.
+We report the detailed results on Monash and ETT in Appendix A.5.2 and A.5.3.
+It can be seen that the results are not that good for PatchTST(ZS) on Monash.
+This is expected since our pretrain data loader will predominantly have context lengths of 512 instead of shorter context lengths as in Monash.
+Moreover the PatchTST model does fewer iterations at the same number of FLOPS.
+On ETT datasets, the PatchTST(ZS) model is performs similarly to TimesFM(ZS) and PatchTST.
+This is also expected since the context length for this study is indeed 512.
+As PatchTST(ZS) is an encoder-decoder model, to pretrain it for zero shot forecasting one should theoretically prepare all possible context lengths and horizon lengths in the pretrain datasets.
+Pretraining it to its maximum performance requires much more compute and likely more careful tuning compared to pretraining TimesFM.
+A.5 Additional Empirical Results
+In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1. The AM based aggregated metrics are presented in Figure 4.
+(a) Monash Archive [9]
+(b) Darts [14]
+(c) ETT (Horizons 96 and 192) [44]
+Figure 4: We report average performance in three groups of datasets.
+In all figures, the lower the metric the better and the error bars represent one standard error.
+Note that among the baselines only TimesFM and llmtime are zero-shot.
+In (a) we report results on the Monash datasets.
+Since the datasets have different scales, we take the Arithmetric Mean (AM) the MAE’s scaled by the MAE of a naive baseline.
+We can see that TimesFM is within significance of the top model N-BEATS.
+In (b), we report the similarly scaled MAE on the Darts benchmarks.
+TimesFM is within significance of the top of method which is ARIMA in this case.
+Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones.
+Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total.
+TimesFM and PatchTST are the best performing models in this case.
+A.5.1 Darts
 We present the MAE results individually from all 8 datasets in Table 3.
 It can be seen that TimesFM performs well for all datasets with clear seasonal patterns.
-We do not perform the among the top-3 only in Sunspot and Wooly datasets.
-On an average we are within significant level of the best model and a close second on the mean scaled MAE metric.
+On an average we are within significant level of the best model.
+Note that there are only 8 time-series as a whole in Darts and theerfore these evaluations have very wide confidence intervals.
+In Figure 8 we present visual comparisons of our forecasts vs some of the baselines.
 Table 3: MAE for Darts datasets. We also include the naive baseline that predicts the last values in the context repeatedly.
-gp arima TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) NAIVE
-AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 14.75 81.45
-AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 10.25 96.35
-GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.69 2.29
-MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 22.46 85.71
-SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 50.88 48.24
-WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2462.11 4075.28
-WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 917.10 1210.33
-HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.44 5.92
-A.2.2 Monash
+GP ARIMA TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) PatchTST NAIVE
+AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 62.51 44.65 81.45
+AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 11.94 21.97 96.35
+GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.50 2.67 2.29
+MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 28.09 42.60 85.71
+SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 41.40 62.33 48.24
+WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2871.33 2498.69 4075.28
+WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 728.92 542.28 1210.33
+HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.85 6.74 5.92
+Scaled MAE (Arithmetic Mean) 0.8193 0.6045 0.8427 0.9176 0.8109 0.6641 0.6829 0.7462 1.0000
+Scaled MAE (Geometric Mean) 0.7509 0.5219 0.7946 0.7316 0.6936 0.4882 0.5767 0.6458 1.0000
+A.5.2 Monash
 Table 4: We present the mean MAE results for our methods along size Monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
-llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
+llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer PatchTST(ZS) TimesFM(ZS) NAIVE
 Dataset
-australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 523.12 659.60
-bitcoin 1.75e18 5.33e18 5.33e18 9.9e18 1.10e18 3.62e18 6.66e18 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.97e18 5.32e18
-pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 58.35 170.88
-weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.17 2.36
-nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 4.09 8.26
-nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 14.33 16.71
-tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 93572.42 99456.05
-tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 10037.84 15845.10
-tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 2916.02 5636.83
-cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 896414.41 386526.37
-covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 555.24 2653.98
-fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 1823.16 2825.67
-traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.03
-traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.13 1.19
-saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 24.53 21.50
-us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 408.49 1152.67
-hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 19.34 24.07
-solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1184.01 1729.41
-Table 3: We present the mean msMAPE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
-llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
-Dataset
-australian electricity demand 15.30 22.07 22.18 13.70 44.22 28.75 8.76 8.35 9.22 12.77 7.74 8.07 8.68 17.27 22.07
-bitcoin 22.70 30.31 40.36 20.12 20.65 31.11 21.48 29.78 21.00 21.16 31.95 22.16 23.32 42.00 192.46
-pedestrian counts 53.02 121.39 122.08 119.76 148.48 138.58 40.29 45.54 39.30 36.10 54.15 34.22 36.16 50.73 121.35
-weather 34.05 50.85 56.19 58.06 51.47 57.98 106.01 59.12 38.17 36.46 50.87 40.13 35.35 50.85 36.81
-nn5 daily 47.43 35.38 21.93 21.11 21.49 25.91 30.20 24.04 23.30 23.75 28.47 22.65 23.18 23.46 48.09
-nn5 weekly 12.28 12.24 11.96 11.62 12.29 11.83 11.45 11.67 11.49 11.52 10.93 14.95 14.83 11.01 13.26
-tourism yearly 48.28 34.10 31.93 33.94 36.52 33.39 46.92 31.54 33.73 34.06 30.24 28.80 34.66 30.14 42.14
-tourism quarterly 25.25 27.41 15.37 17.16 15.07 16.58 15.86 16.53 16.20 15.29 14.45 15.56 16.97 19.92 31.68
-tourism monthly 33.53 36.39 19.89 21.20 19.02 19.73 21.11 21.10 20.11 18.35 20.42 18.92 19.74 22.77 40.40
-cif 2016 18.63 14.94 13.04 12.19 12.18 11.69 12.32 14.86 12.31 13.54 11.71 18.82 12.55 20.00 15.41
-covid deaths 20.65 15.35 15.57 8.71 8.64 9.26 18.34 15.40 18.58 34.18 32.36 14.50 40.71 15.89 134.79
-fred md 9.75 8.72 9.72 7.97 8.40 7.98 30.77 9.16 8.99 8.33 8.32 9.08 11.26 8.47 8.76
-traffic hourly 11.70 8.73 8.73 12.58 9.84 11.72 5.97 7.94 4.30 4.16 5.16 5.22 4.10 4.76 8.73
-traffic weekly 12.85 12.40 12.48 12.80 12.63 13.45 12.46 12.90 12.64 13.13 12.31 13.21 15.18 12.48 12.98
-saugeenday 58.53 35.99 35.97 37.34 67.50 37.55 45.32 35.55 39.32 40.22 56.03 37.02 56.62 43.31 35.99
-us births 4.48 11.77 5.82 3.81 4.05 5.17 5.75 4.23 5.55 4.13 4.17 4.88 4.36 3.94 11.35
-hospital 21.68 17.94 17.27 17.55 17.46 17.79 17.56 18.04 18.29 17.41 17.72 17.51 20.04 17.86 21.55
-solar weekly 37.33 24.59 24.76 19.05 22.93 17.87 21.65 29.35 21.52 15.00 24.05 32.50 12.26 24.20 32.80
+australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 382.23 448.81 659.60
+bitcoin 1.75e18 5.33e18 5.33e18 9.90e17 1.10e18 3.62e18 6.66e17 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.11e18 1.3e18 7.77e17
+pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 51.27 40.71 170.88
+weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.07 2.07 2.36
+nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 3.77 3.54 8.26
+nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 17.00 14.67 16.71
+tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 224411.89 109977.29 99456.05
+tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 21276.98 12102.04 15845.10
+tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 4596.21 3183.77 5636.83
+cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 8374813.14 773980.44 386526.37
+covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 348.60 209.80 353.71
+fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 4965.51 947.12 2825.67
+traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.01 0.03
+traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.23 1.12 1.19
+saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 22.33 24.63 21.50
+us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 1193.28 437.27 1152.67
+hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 20.87 19.41 24.07
+solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1093.46 1258.27 1729.41
+Scaled MAE (Arithmetic Mean) 1.0588 1.3115 1.2295 0.8824 0.9337 1.2427 1.0382 0.8924 0.8942 1.1925 0.7844 1.8119 1.4840 2.1419 0.8005 1.0000
+Scaled MAE (Geometric Mean) 0.9715 1.0855 0.9371 0.7736 0.8104 0.9449 0.8218 0.7733 0.7044 0.7477 0.7005 0.9384 0.8619 1.0557 0.6846 1.0000
 In Table 4 we present the actual MAE numbers that are behind the main Figure 2(a).
-It should be noted that though the scaled MAE metric accounts for scale variation across datasets, it does not account for within-dataset scale variations.
-Indeed, in datasets like cif 2016 and covid deaths, there are time-series with different order of magnitudes.
-The mean MAE for these datasets does not reflect average performance but only the performance of the higher scale time-series.
-Therefore in Table 3 we also present the normalized msMAPE metrics for TimesFM and all baselines.
-A.2.3 Informer
+In Figure 9, we present some examples of our zero-shot forecasts.
+For most datasets, we set the context window to be the maximum length of the series in the dataset capped at 512 (similar to statistcal models used in the official Monash baselines).
+For some datasets, we did some inference time-tuning of the context length, i.e we predict the last horizon length number of points in the training set with context lengths 32, 64 and maximum allowed and chose the best one in terms of this validation metric.
+This is fair as most Monash DL baselines use different context lengths for different datasets during training and our model is completely zero-shot.
+The max context lengths used for these datasets are (cif 2016, 32), (tourism yearly, 32), (covid deaths, 32), (bitcoin, 32), (tourism monthly, 32) and (tourism monthly, 64).
+A.5.3 Informer
 We present the MAE on the last split of the test set for all dataset, horizon pairs considered in Table 5.
@@
 Table 5: MAE for ETT datasets for prediction horizons 96 and 192. Owing to expensive evaluations for llmtime, the results are reported on the last test window of the original test split.
-llmtime(ZS) PatchTST TimesFM(ZS) FEDFormer AutoFormer Informer
+llmtime(ZS)* PatchTST PatchTST(ZS) FEDFormer AutoFormer Informer TimesFM(ZS)
 Dataset
-ETTh1 (Horizon=96) 0.42 0.41 0.37 0.58 0.55 0.76
-ETTh1 (Horizon=192) 0.50 0.49 0.49 0.64 0.64 0.78
-ETTh2 (Horizon=96) 0.33 0.28 0.28 0.67 0.65 1.94
-ETTh2 (Horizon=192) 0.70 0.68 0.58 0.82 0.82 2.02
-ETTm1 (Horizon=96) 0.37 0.33 0.25 0.41 0.54 0.71
-ETTm1 (Horizon=192) 0.71 0.31 0.24 0.49 0.46 0.68
-ETTm2 (Horizon=96) 0.29 0.23 0.28 0.36 0.29 0.48
-ETTm2 (Horizon=192) 0.31 0.25 0.24 0.25 0.30 0.51
-A.3 More Details on Models
+ETTh1 (horizon=96) 0.42 0.41 0.39 0.58 0.55 0.76 0.45
+ETTh1 (horizon=192) 0.50 0.49 0.50 0.64 0.64 0.78 0.53
+ETTh2 (horizon=96) 0.33 0.28 0.37 0.67 0.65 1.94 0.35
+ETTh2 (horizon=192) 0.70 0.68 0.59 0.82 0.82 2.02 0.62
+ETTm1 (horizon=96) 0.37 0.33 0.24 0.41 0.54 0.71 0.19
+ETTm1 (horizon=192) 0.71 0.31 0.26 0.49 0.46 0.68 0.26
+ETTm2 (horizon=96) 0.29 0.23 0.22 0.36 0.29 0.48 0.24
+ETTm2 (horizon=192) 0.31 0.25 0.22 0.25 0.30 0.51 0.27
+Avg 0.45 0.37 0.35 0.53 0.53 0.99 0.36
+A.6 More Details on Models
 We now present implementation details about TimesFM and other baselines.
+TimesFM.
+For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128.
+The model dimension is set to 1280.
+We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $45e-$ 4.
+The hyper-parameters of TimesFM for various sizes are provided in Table 6.
+Note that the settings are for the base models and not ablation models.
+The hidden dims of both the residual block and the FFN in the transformer layers are set as the same as model dimensions.
+We keep layer norm in transformer layers but not in the residual blocks.
+Table 6: Hyper-parameters for TimesFM
+num_layers model_dims output_patch_len input_patch_len num_heads dropout
+Size
+200M 20 1280 128 32 16 0.2
+70M 10 1024 128 32 16 0.2
+17M 10 512 128 32 16 0.2
 Monash Baselines.
@@
 Darts Baselines. For all the Darts baselines we use the precomputed outputs provided by the authors of  [11]. For more details please see Section C.1 in that paper.
-TimesFM.
-For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128.
-The model dimension is set to 1280.
-We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $45e-$ 4.
 Informer Baselines.
@@
 Note that as of January 2024, OpenAI has discontinued access to GPT-3, therefore we had to use the GPT-3.5-Turbo model.
-A.4 Synthetic Data
-We create the synthetic data to reflect common time series patterns using traditional statistical models. We start with four simple times series patterns:
+A.7 Date Features
+As we mentioned earlier, since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time.
+However, the datetime column is ubiquitous in all time-series data, so we can technically have date derived features like day of the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$.
+If so, the learning task can be rewritten as
+$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L},\mathbf{x}_{1:L+H}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$
+There are many options to incorporate these features into the model, one being to directly concatenate them after the time-points in each patch.
+For this paper we decide to focus on the univariate time-series input, and will investigate this enhancement in the future.
+A.8 Synthetic Data
+We create the synthetic data to reflect common time-series patterns using traditional statistical models. We start with four simple times series patterns:
   • Piece-wise linear trends (I), where the number of the piece-wise linear components is randomly chosen between 2 and 8.
   • ARMA$(p,q)$ (II), where $1\leq p,q\leq 8$ and the corresponding coefficients are generated from either a multivariate Gaussian or a uniform, then normalized.
-  • Seasonal patterns. In particular we create the sine (III) and the cosine (IV) waves of different random periods between 4 and 96 time-points and time delays.
-We then randomly enable / disable these four components (I) - (IV), generate their time series of length 1024 respectively, and sum them up using uniformly sampled random weights to create each times series in the synthetic datasets.
+  • Seasonal patterns. In particular we create the sine (III) and the cosine (IV) waves of different random periods between 4 and max context length / 2 time-points and time delays.
+We then randomly enable / disable these four components (I) - (IV), generate their time-series of length 2048 respectively, and sum them up using uniformly sampled random weights to create each times series in the synthetic datasets.
 We also choose to apply the trend multiplicatively 50% of the times the trend component is chosen.
+A.9 Illustrative Examples
+We conduct a visual inspection of the forecasts generated by TimesFM, first on some synthetic examples and then on the benchmark datasets.
+In Figure 5 we show 4 different synthetic curves: (1) sum of 5 sine curves of different periods, (2) a sine curve linearly scaled, (3) a sine curve with a linear trend, and (4) minimum of two sine curves with a linear trend.
+Our results suggests that TimesFM picks up the trend and seasonal components readily interpretable by humans, while ARIMA and (to a lesser extent) llmtime fail in some of the instances.
+As illustrated in Figure 6, TimesFM also effectively captures these subtle characteristics within both the trend and seasonal patterns of the depicted real world time-series.
+For instance, in the Air Passenger dataset, TimesFM correctly captures the amplitude increase with trend –this is also reflected by the fact that it attains the best MAE on this dataset (see Table 3).
+In the traffic hourly example on the left, it can be seen that TimesFM can correctly identify the seasonal peaks even in the presence of outliers in the context, while llmtime is thrown off.
+Figure 5: Forecasts visualized on synthetic curves. The bottom row plots zoom in on the prediction horizon for the sake of clarity.
+Figure 6: Forecasts visualized on Darts and Monash. The bottom row plots zoom in on the prediction horizon for the sake of clarity.
+We provide more visualization in Figure 7, Figure 8 and Figure 9.
+Figure 7: Forecasts visualized on synthetic curves. The second row plots zoom in on the prediction horizon for the sake of clarity.
+Figure 8: Forecasts visualized on all Darts datasets. The second row plots zoom in on the prediction horizon for the sake of clarity.
+Figure 9: Forecasts visualized on a few Monash datasets. The second row plots zoom in on the prediction horizon for the sake of clarity.